### v7

# Integral Invariants and Hamiltonian Systems

Abstract.
In this review and methodological article we discuss the main ideas of the integral invariants theory. This theory was originated by Poincare and Cartan, [7], [2]. We show how ideas of this theory connect such different fields of mathematical physics as Hamiltonian dynamics, optics and hydrodynamics. We focus our attention on the results that are rarely expounded in the textbooks.
1. Invariant Differential Forms
Consider a smooth dynamical system
$\dot{x}=v(x)$ (1)
defined on the smooth manifold $M$ with local coordinates $x=(x^{1},\ldots,x^{m})^{T}.$
Remark 1.
Throughout all this text we assume all the objects to be as smooth as the formulas require to make sense and all the integrals are assumed to be correctly defined. When we use the Stokes theorem we assume that the corresponding conditions are satisfied etc.
We use
$g^{t}(x),\quad g^{t}:M\to M$
to denote the phase flow of this system:
$\frac{d}{dt}g^{t}(\hat{x})=v(g^{t}(\hat{x})),\quad g^{0}(\hat{x})=\hat{x}$.
Recall that $g^{t+s}=g^{t}\circ g^{s}.$
We use the following trivial formula:
$g^{t}(x)=x+v(x)t+o(t),\quad t\to 0.$ (2)
Definition 1 ([6]).
The following differential operator
$L_{v}\omega=\frac{d}{dt}\Big|_{t=0}g^{t}_{*}\omega=\frac{d}{dt}\Big|_{t=0}\sum_{i_{1}<\ldots<i_{k}}\omega_{i_{1}\ldots i_{k}}\big(g^{t}(x)\big)d\Big(\big(g^{t}(x)\big)^{i_{1}}\Big)\wedge\ldots\wedge d\Big(\big(g^{t}(x)\big)^{i_{k}}\Big)$
is called the Lie derivative of the differential form
$\omega=\sum_{i_{1}<\ldots<i_{k}}\omega_{i_{1}\ldots i_{k}}(x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}}.$
Recall the Cartan homotopy formula:
$L_{v}\omega=di_{v}\omega+i_{v}d\omega.$
The operator
$i_{v}\omega=\sum_{i_{2}<\ldots<i_{k}}\omega_{ii_{2}\ldots i_{k}}v^{i}dx^{i_{2}}\wedge\ldots\wedge dx^{i_{k}}$
is called the interior derivative.
This operator possesses the following property:
$i_{v}(\omega\wedge\nu)=(i_{v}\omega)\wedge\nu+(-1)^{k}\omega\wedge(i_{v}\nu),$ (3)
where $k$ is degree of $\omega$.
Recall that differentiation has the same property:
$d(\omega\wedge\nu)=(d\omega)\wedge\nu+(-1)^{k}\omega\wedge(d\nu).$
Problem 1.
By using (2) show that
$L_{v}\big(\omega_{i}dx^{i}\big)$ $=\Big(\frac{\partial\omega_{i}}{\partial x^{s}}v^{s}+\frac{\partial v^{s}}{\partial x^{i}}\omega_{s}\Big)dx^{i},$ (4)
$L_{v}\Big(\rho(x)dx^{1}\wedge\ldots\wedge dx^{m}\Big)$ $=\frac{\partial(\rho v^{i})}{\partial x^{i}}dx^{1}\wedge\ldots\wedge dx^{m},$
$L_{v}f$ $=v^{i}\frac{\partial f}{\partial x^{i}},\quad f:M\to\mathbb{R},$
$L_{v}d$ $=dL_{v}.$ (5)
Definition 2.
A $k$-form $\omega$ is called the integral invariant of system (1) iff $L_{v}\omega=0$.
A $k$-form $\omega$ is called the relative integral invariant of system (1) iff there exists a k1k-1 form $\Omega$ such that $L_{v}\omega=d\Omega$.
Remark 2.
From formula (5) it follows that if the form $\omega$ is a relative integral invariant then the form $d\omega$ is an integral invariant.
Problem 2.
Show that if the integral invariant $\omega$ is a closed form:
$L_{v}\omega=0,\quad d\omega=0,$
and the form $\psi$ is a relative integral invariant:
$L_{v}\psi=d\Psi$
Then $\omega\wedge\psi$ is a relative integral invariant:
$L_{v}(\omega\wedge\psi)=(-1)^{\mathrm{deg}\,\omega}d(\omega\wedge\Psi).$
Let $\Sigma\subset M$ stand for a smooth $k$-dimensional manifold that is smoothly embedded in $M$.
Theorem 1.
The following formula holds
$\frac{d}{dt}\Big|_{t=0}\int_{g^{t}(\Sigma)}\omega=\int_{\Sigma}L_{v}\omega.$ (6)
Proof
We prove this formula for the case when $\Sigma$ is covered with a single chart. We use
$u:D\to M,\quad u(D)=\Sigma$
to denote the corresponding embedding. Here $D\subset\mathbb{R}^{k}$ is an open domain.
Thus we have:
$\int_{g^{t}(\Sigma)}\omega=\int_{D}(g^{t}\circ u)_{*}\omega=\int_{D}(u_{*}\circ(g^{t})_{*})\omega,$
and
$\frac{d}{dt}\Big|_{t=0}\int_{D}(u_{*}\circ(g^{t})_{*})\omega=\int_{D}u_{*}\frac{d}{dt}\Big|_{t=0}(g^{t})_{*}\omega=\int_{D}u_{*}(L_{v}\omega).$
The theorem is proved.
Corollary 1.
Formula (6) implies
$\frac{d}{dt}\int_{g^{t}(\Sigma)}\omega=\frac{d}{ds}\Big|_{s=0}\int_{g^{t+s}(\Sigma)}\omega=\frac{d}{ds}\Big|_{s=0}\int_{g^{s}\big(g^{t}(\Sigma)\big)}\omega=\int_{g^{t}(\Sigma)}L_{v}\omega.$ (7)
Theorem 2.
If $\omega$ is an integral invariant then for any admissible (see Remark 1) $k$-dimensional submanifold $\Sigma\subset M$ the following integral
$\int_{g^{t}(\Sigma)}\omega$
does not depend on $t$. The converse is also true.
Indeed, this follows from (7) directly.
Theorem 3.
Assume that $\omega$ is a relative integral invariant. Then for any compact $k$-dimensional submanifold $\Sigma\subset M,\quad\partial\Sigma=\emptyset$ the following integral
$\int_{g^{t}(\Sigma)}\omega$
does not depend on $t$.
Indeed, from formula (7) and by the Stokes theorem we have
$\frac{d}{dt}\int_{g^{t}(\Sigma)}\omega=\int_{g^{t}(\Sigma)}L_{v}\omega=\int_{g^{t}(\Sigma)}d\Omega=0.$
Here we take into account that $\partial g^{t}(\Sigma)=\emptyset$.
2. Invariant Differential Forms of the Systems with First Integrals
2.1. The Case $\dim M=2$
If $m=2$ and $q$ is an invariant $2$-form of system (1):
$L_{v}q=0,$
then $i_{v}q$ is a closed form. Indeed:
$L_{v}q=di_{v}q+i_{v}dq,\quad dq=0.$
Thus locally there exists a function $f$ such that
$df=i_{v}q;$
and
$L_{v}f=i_{v}df=i_{v}i_{v}q=0.$
Consequently the function $f$ is a first integral.
2.2. Invariant Measure on First Integral Level
Assume that system (1) has a first integral $F$:
$L_{v}F=0,\quad dF\neq 0$
and an invariant $m$-form $\omega:\quad L_{v}\omega=0$. Without loss of generality we can locally assume that
$\frac{\partial F}{\partial x^{m}}\neq 0.$
Then
$\omega=\rho(x)dx^{1}\wedge\ldots,dx^{m}=\rho(x)\Big(\frac{\partial F}{\partial x^{m}}\Big)^{-1}dx^{1}\wedge\ldots dx^{m-1}\wedge dF.$
In the invariant terms this means that the form $\omega$ is presented as the following exterior product
$\omega=\lambda\wedge dF,$
where $\lambda$ is an m1m-1-form which is defined up to an additive form $\gamma,\quad\gamma\wedge dF=0$.
Proposition 1.
The following restricted differential form
$\lambda\mid_{Z},\quad Z=\{F=const\}$
is an integral invariant for the restriction of system (1) to the manifold $Z$.
Indeed,
$L_{v}\omega=(L_{v}\lambda)\wedge dF+\lambda\wedge(L_{v}dF),\quad L_{v}(dF)=dL_{v}F=0.$
Therefore we have $(L_{v}\lambda)\wedge dF=0.$ Let $e_{1},\ldots,e_{m}$ be a basis in $T_{x}M$ such that $e_{1},\ldots,e_{m-1}$ is a basis in $T_{x}Z$. We consequently obtain
$\big((L_{v}\lambda)\wedge dF\big)(e_{1},\ldots,e_{m})=c(L_{v}\lambda)(e_{1},\ldots,e_{m-1})\cdot dF(e_{m})=0.$
Problem 3.
Calculate the constant $c$ and make sure that $c\neq 0$.
[…]
The Proposition is proved.
Taking into account the results of section 2.1 we see that if system (1) has m2m-2 independent first integrals and an invariant $m$-form:
$L_{v}\omega=0,\quad\omega=\rho(x)dx^{1}\wedge\ldots,dx^{m},\quad\rho(x)>0$
then this system is integrated in the closed form.
3. The Lie Derivative in the Nonautonomous case: the Helmholtz and Kelvin Theorems
3.1. General Construction
Consider a vector field
$v=(v^{1},\ldots v^{m})(t,x),\quad(t,x)\in\tilde{M}=(t_{1},t_{2})\times M.$
We use $G^{t}_{t_{0}}:M\to M$ to denote the shift along the trajectories of the system
$\dot{x}=v(t,x),\quad\frac{dG^{t}_{t_{0}}(x)}{dt}=v(t,G^{t}_{t_{0}}(x)),\quad G^{t_{0}}_{t_{0}}(x)=x,\quad G_{t_{1}}^{t}\circ G_{t_{0}}^{t_{1}}=G_{t_{0}}^{t}.$ (8)
Let $g^{\tau}:\tilde{M}\to\tilde{M}$ be a flow of the system
$\frac{dz}{d\tau}=\tilde{v}(z),\quad z=(t,x^{1},\ldots,x^{m})^{T},\quad\tilde{v}=(1,v^{1},\ldots,v^{m})^{T}.$ (9)
Here $\tilde{M}$ denotes the extended phase space of (8).
The following equality holds
$g^{\tau}(t_{0},x)=(t_{0}+\tau,G_{t_{0}}^{t_{0}+\tau}(x)).$
Remark 3.
System (9) has an integral invariant $dt.$ Indeed,
$L_{\tilde{v}}(dt)=dL_{\tilde{v}}(t)=d\Big(1\cdot\frac{\partial t}{\partial t}\Big)=0.$
Introduce a form
$\omega=\sum_{i_{1}<\ldots<i_{k}}\omega_{i_{1}\ldots i_{k}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}}$
and fix the local coordinates.
We can consider the form $\omega$ as a form on the manifold $\tilde{M}$ or as a form on the manifold $M$. In the last case $t$ is a parameter.
Introduce the notations
$\frac{\partial\omega}{\partial t}=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}(t,x)}{\partial t}dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}},\quad\frac{\partial v}{\partial t}=\Big(\frac{\partial v^{1}}{\partial t},\ldots,\frac{\partial v^{m}}{\partial t}\Big);$
and
$d$ $=d_{t}+d_{x},\quad d_{t}\omega=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}(t,x)}{\partial t}dt\wedge dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}};$
$d_{x}\omega$ $=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}}{\partial x^{l}}dx^{l}\wedge dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}},$
$L_{v}\omega$ $:=d_{x}i_{v}\omega+i_{v}d_{x}\omega.$
For instance we have $d_{t}\omega=(dt)\wedge\frac{\partial\omega}{\partial t}$.
Theorem 4.
The following formula holds
$L_{\tilde{v}}\omega=\frac{\partial\omega}{\partial t}+L_{v}\omega+(dt)\wedge i_{\frac{\partial v}{\partial t}}\omega.$
Proof of Theorem 4
Let us split the vector field into the following summands
$\tilde{v}=e+v_{*},\quad e=(1,0,\ldots,0),\quad v_{*}=(0,v^{1},\ldots,v^{m}).$
Correspondingly $L_{\tilde{v}}=L_{e}+L_{v_{*}}.$
Observe that
$L_{e}\omega=\frac{\partial\omega}{\partial t},\quad i_{v_{*}}\omega=i_{v}\omega.$
By the homotopy formula we get
$L_{v_{*}}\omega=i_{v_{*}}d_{x}\omega+i_{v_{*}}d_{t}\omega+d_{x}i_{v_{*}}\omega+d_{t}i_{v_{*}}\omega.$ (10)
Since
$i_{v_{*}}d_{x}\omega=i_{v}d_{x}\omega,\quad d_{x}i_{v_{*}}\omega=d_{x}i_{v}\omega,$
the equality (10) takes the form
$L_{v_{*}}\omega=L_{v}\omega+i_{v_{*}}d_{t}\omega+d_{t}i_{v_{*}}\omega.$
By formula (3) it follows that:
$i_{v_{*}}d_{t}\omega=i_{v_{*}}\Big((dt)\wedge\frac{\partial\omega}{\partial t}\Big)=-(dt)\wedge i_{v}\frac{\partial\omega}{\partial t},$
and
$d_{t}i_{v_{*}}\omega=(dt)\wedge\frac{\partial i_{v}\omega}{\partial t}=(dt)\wedge\Big(i_{\frac{\partial v}{\partial t}}\omega+i_{v}\frac{\partial\omega}{\partial t}\Big).$
This proves the theorem.
Theorem 5.
For any admissible $k$-dimensional submanifold $A\subset M$ one has
$\frac{d}{dt}\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{G^{t}_{t_{0}}(A)}\Big(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\Big).$ (11)
Proof
Introduce a manifold
$A_{t}=\{t\}\times A\subset\tilde{M},\quad g^{\tau}(A_{t})=\{t+\tau\}\times G^{t+\tau}_{t}(A)\subset\tilde{M}$.
Using theorem 4 and formula 7 calculate:
$\frac{d}{d\tau}\int_{g^{\tau}(A_{t_{0}})}\omega$ $=\int_{g^{\tau}(A_{t_{0}})}L_{\tilde{v}}\omega=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}(L_{\tilde{v}}\omega)\mid_{t=t_{0}+\tau}$
$=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}\Big(\frac{\partial\omega}{\partial t}+L_{v}\omega\Big)\Big|_{t=t_{0}+\tau}.$
To finish the proof it remains to observe that
$\int_{g^{\tau}(A_{t_{0}})}\omega=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}\omega\mid_{t=t_{0}+\tau}.$
The theorem is proved.
Remark 4.
After the change of variables
$x\mapsto y=G^{t}_{t_{0}}(x)$
formula (11) takes the form
$\frac{d}{dt}\int_{A}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\int_{A}\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\int_{A}(G^{t}_{t_{0}})_{*}\Big(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\Big).$
Since AA is an arbitrary set we have
$\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=(G^{t}_{t_{0}})_{*}\Big(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\Big).$ (12)
Theorem 6.
1) Let $A\subset M$ be an admissible $k$-dimensional submanifold and
$\frac{\partial\omega}{\partial t}+L_{v}\omega=0.$
Then we have
$\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{A}\omega(t_{0},\cdot).$
2) Suppose that there exists a k1k-1-form
$\Omega=\sum_{i_{1}<\ldots<i_{k-1}}\Omega_{i_{1}\ldots i_{k-1}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k-1}}$
such that
$\frac{\partial\omega}{\partial t}+L_{v}\omega=d_{x}\Omega$
and $A\subset M$ is a compact submanifold without boundary: $\partial A=\emptyset$. Then we have
$\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{A}\omega(t_{0},\cdot).$
Theorem 6 follows from theorem 5 in the same manner as theorems 3, 2 follow from formula (7).
The theorem is proved.
Problem 4.
Prove that if
$L_{v}\omega+\frac{\partial\omega}{\partial t}=d_{x}\Omega,$
then
$L_{\tilde{v}}\big(dt\wedge\omega\big)=-d\big(dt\wedge\Omega\big).$
I.e. in this situation the form $dt\wedge\omega$ is a relative integral invariant of system (9). In particular if $\Omega=0$ then $dt\wedge\omega$ is an integral invariant of (9).
3.2. Applications to Hydrodynamics
In this section we put $M=\mathbb{R}^{3}$ and let $x^{1},x^{2},x^{3}$ be the standard right-hand oriented Euclidean frame.
In this section we also put $d=d_{x}$.
Let
$\boldsymbol{A}(t,x)=(A_{i}),\quad\boldsymbol{B}(t,x)=(B_{i})$
be vector fields in $\mathbb{R}^{3}$ and let $f(t,x)$ be a scalar function. There are correspondences
$f\mapsto\omega^{3}_{f}=fdx^{1}\wedge dx^{2}\wedge dx^{3}$
and
$\boldsymbol{A}\mapsto\omega^{1}_{\boldsymbol{A}}=A_{i}dx^{i},\quad\boldsymbol{A}\mapsto\omega^{2}_{\boldsymbol{A}}=A_{1}dx^{2}\wedge dx^{3}+A_{2}dx^{3}\wedge dx^{1}+A_{3}dx^{1}\wedge dx^{2}.$
These correspondences can be expressed in terms of the Hodge star operation.
The following formulas are checked by direct calculation:
$df=\omega^{1}_{\mathrm{grad}\,f},\quad d\omega^{1}_{\boldsymbol{A}}=\omega^{2}_{\mathrm{curl}\,\boldsymbol{A}},\quad d\omega^{2}_{\boldsymbol{A}}=\omega^{3}_{\mathrm{div}\,\boldsymbol{A}}$
and
$i_{\boldsymbol{B}}\omega^{1}_{\boldsymbol{A}}=(\boldsymbol{A},\boldsymbol{B}),\quad i_{\boldsymbol{B}}\omega^{2}_{\boldsymbol{A}}=\omega^{1}_{\boldsymbol{A}\times\boldsymbol{B}},\quad i_{\boldsymbol{B}}\omega^{3}_{f}=f\omega^{2}_{\boldsymbol{B}}.$
These formulas imply the next theorem.
Theorem 7.
The following formulas hold
$\frac{\partial\omega^{1}_{\boldsymbol{A}}}{\partial t}+L_{v}\omega^{1}_{\boldsymbol{A}}$ $=\omega^{1}_{\frac{\partial\boldsymbol{A}}{\partial t}+(\mathrm{curl}\,\boldsymbol{A})\times v}+d(v,\boldsymbol{A}),$
$\frac{\partial\omega^{2}_{\boldsymbol{A}}}{\partial t}+L_{v}\omega^{2}_{\boldsymbol{A}}$ $=\omega^{2}_{\frac{\partial\boldsymbol{A}}{\partial t}+\mathrm{curl}\,(\boldsymbol{A}\times v)+v\,\mathrm{div}\,\boldsymbol{A}},$
$\frac{\partial\omega^{3}_{f}}{\partial t}+L_{v}\omega^{3}_{f}$ $=\omega^{3}_{\frac{\partial f}{\partial t}+\mathrm{div}\,(fv)}.$
Theorem 8.
Let
$\gamma,\Sigma,D\subset\mathbb{R}^{3}$
be a closed curve, a two-dimensional surface and a domain respectively.
1) If
$\frac{\partial\boldsymbol{A}}{\partial t}+(\mathrm{curl}\,\boldsymbol{A})\times v=\mathrm{grad}\,\psi,\quad\psi=\psi(t,x),$ (13)
then
$\int_{G_{t_{0}}^{t}(\gamma)}\omega^{1}_{\boldsymbol{A}(t,\cdot)}=\int_{\gamma}\omega^{1}_{\boldsymbol{A}(t_{0},\cdot)};$
2) if
$\frac{\partial\boldsymbol{A}}{\partial t}+\mathrm{curl}\,(\boldsymbol{A}\times v)+v\,\mathrm{div}\,\boldsymbol{A}=0,$ (14)
then
$\int_{G_{t_{0}}^{t}(\Sigma)}\omega^{2}_{\boldsymbol{A}(t,\cdot)}=\int_{\Sigma}\omega^{2}_{\boldsymbol{A}(t_{0},\cdot)};$
3) if
$\frac{\partial f}{\partial t}+\mathrm{div}\,(fv)=0,$ (15)
then
$\int_{G_{t_{0}}^{t}(D)}\omega^{3}_{f(t,\cdot)}=\int_{D}\omega^{3}_{f(t_{0},\cdot)}.$
This theorem is a consequence from theorems 6 and 7.
Problem 5.
Prove the equality
$\mathrm{curl}\,(\boldsymbol{A}\times\boldsymbol{B})=[\boldsymbol{A},\boldsymbol{B}]+\boldsymbol{A}\,\mathrm{div}\,\boldsymbol{B}-\boldsymbol{B}\,\mathrm{div}\,\boldsymbol{A}.$
We use square brackets to denote the commutator of the vector fields:
$\boldsymbol{A},\boldsymbol{B}]=\frac{\partial\boldsymbol{A}}{\partial x}\boldsymbol{B}-\frac{\partial\boldsymbol{B}}{\partial x}\boldsymbol{A}.$
Hint: prove and use the formula
$L_{u}i_{v}-i_{v}L_{u}=i_{[u,v]}.$
Let $v(t,x)$ be a velocity field of the ideal fluid which undergoes potential forces and such that its density depends on pressure only.
Under these assumptions equation (14) holds with $\boldsymbol{A}=\mathrm{curl}\,v$. The velocity field $v$ satisfies equation (13) with $\boldsymbol{A}=v$ as well [3].
The famous Helmholtz and Kelvin theorems follow from theorem 8.
Moreover (15) is the continuity equation.
3.3. Other Remarks
Proposition 2.
The following formula holds
$L_{\tilde{v}}\big((dt)\wedge\omega\big)=(dt)\wedge\Big(\frac{\partial\omega}{\partial t}+L_{v}\omega\Big).$
Indeed,
$L_{\tilde{v}}\big((dt)\wedge\omega\big)=\big(L_{\tilde{v}}(dt)\big)\wedge\omega+(dt)\wedge L_{\tilde{v}}\omega,$
and $L_{\tilde{v}}(dt)=dL_{\tilde{v}}t=0.$ Now the proposition follows from theorem 4.
Any $k$-form in $\tilde{M}$ can be presented in the shape
$\mu=\mu^{+}+dt\wedge\mu^{-},$
where
$\mu^{+}$ $=\sum_{i_{1}<\ldots<i_{k}}\mu^{+}_{i_{1}\ldots i_{k}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}},$
$\mu^{-}$ $=\sum_{i_{1}<\ldots<i_{k-1}}\mu^{-}_{i_{1}\ldots i_{k-1}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k-1}}.$
Combining proposition 2 and theorem 4 we yield the following theorem.
Theorem 9.
The equality holds:
$L_{\tilde{v}}\mu=\frac{\partial\mu^{+}}{\partial t}+L_{v}\mu^{+}+(dt)\wedge\Big(i_{\frac{\partial v}{\partial t}}\mu^{+}+\frac{\partial\mu^{-}}{\partial t}+L_{v}\mu^{-}\Big).$
Formula 12 prompts a way of solving the following Cauchy problem:
$\frac{\partial\omega}{\partial t}(t,x)+L_{v(t,x)}\omega(t,x)=0,\quad\omega\mid_{t=t_{0}}=\hat{\omega}$.
Indeed, from this formula we have
$\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=0.$
Since $(G^{t}_{t_{0}})_{*}\omega(t,\cdot)$ does not depend on $t$ we obtain
$(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\hat{\omega},\quad\omega(t,x)=\big((G^{t}_{t_{0}})^{-1}\big)_{*}\hat{\omega}$.
4. Darboux Theorem [4], [8]
In this section we consider some nontrivial application of the theory built above.
Let $D\subset\mathbb{R}^{2m}=\{x=(x^{1},\ldots,x^{2m})\}$ be an open neighbourhood of the origin.
A differential form
$\omega=\sum_{i<j}\omega_{ij}(x)dx^{i}\wedge dx^{j}$
defined in $D$ is closed and non degenerate:
$d\omega=0,\quad\det(\omega_{ij}(x))\neq 0,\quad x\in D.$
Theorem 10.
In some open neighbourhood $D^{\prime}$ of the origin there are local coordinates $y=(y^{1},\ldots,y^{2m})$ such that the form $\omega$ has constant coefficients:
$\omega=\sum_{i<j}\omega^{\prime}_{ij}dy^{i}\wedge dy^{j},\quad\omega^{\prime}_{ij}=\mathrm{const}_{ij}.$
From vector algebra we know that there is a linear transformation $y\mapsto(q^{1},\ldots,q^{m},p_{1},\ldots,p_{m})$ such that
$\omega=dp_{i}\wedge dq^{i}.$
Proof of Theorem 10
Introduce a form
$\omega_{1}=\omega\mid_{x=0}=\sum_{i<j}\omega_{ij}(0)dx^{i}\wedge dx^{j},\quad d\omega_{1}=0.$
Construct a family of differential forms as follows
$\Omega(t,x)=t\omega_{1}+(1-t)\omega=\sum_{i<j}\Omega_{ij}(t,x)dx^{i}\wedge dx^{j},\quad t\in[0,1].$
These forms are defined in $D$.
Each form $\Omega(t,0)=\omega\mid_{x=0}$ is non degenerate so all the forms $\{\Omega(t,x)\}$ are non degenerate in some neighbourhood of the origin. Since $d(\omega_{1}-\omega)=0$ we have a form
$\alpha=\alpha_{k}(x)dx^{k},\quad\omega_{1}-\omega=d\alpha.$
Let us choose $\alpha$ such that $\alpha\mid_{x=0}=0.$
Define a vector field $v(t,x)$ as follows
$i_{v(t,x)}\Omega(t,x)=-\alpha,\quad v^{i}(t,x)=-\Omega^{ij}(t,x)\alpha_{j}(x).$
Since $v(t,0)=0$ the corresponding family of shifts $x\mapsto G^{t}_{0}(x)$ is defined for small $|x|$ and for $t\in[0,1]$.
From (12) we obtain
$\frac{d}{dt}(G^{t}_{0})_{*}\Omega(t,\cdot)=(G^{t}_{0})_{*}\Big(\frac{\partial\Omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\Omega(t,\cdot)\Big).$
From the Cartan formula we get
$L_{v(t,\cdot)}\Omega(t,\cdot)=i_{v(t,x)}d_{x}\Omega(t,x)+d_{x}i_{v(t,x)}\Omega(t,x)=-d\alpha.$
On the other hand
$\frac{\partial\Omega}{\partial t}=\omega_{1}-\omega=d\alpha.$
Thus we yield
$\frac{d}{dt}(G^{t}_{0})_{*}\Omega(t,\cdot)=0.$
Consequently $(G^{1}_{0})_{*}\Omega(1,\cdot)=(G^{0}_{0})_{*}\Omega(0,\cdot)=\omega$ and the mapping $G^{1}_{0}$ takes the form $\omega_{1}$ to the form $\omega$.
The theorem is proved.
5. Integral Invariants of the Hamilton Equations
Recall that a manifold $N,\quad\dim N=2m$ is called the symplectic manifold if it is endowed with a nondegenerate closed 2-form. Denote this form by $\beta$. Nondegenerateness means that the equality $i_{a}\beta=0$ implies $a=0$.
Darboux’s theorem (see section 4) says that any point of $N$ belongs to a chart with local coordinates
$z=(x^{1},\ldots,x^{m},p_{1},\ldots,p_{m})$
such that
$\beta=dp_{i}\wedge dx^{i}.$
These coordinates are called symplectic or canonical.
Consider a Hamiltonian system with a Hamiltonian function $H=H(t,x,p)$:
$\dot{x}^{i}=\frac{\partial H}{\partial p_{i}},\quad\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}},\quad t\in I=(t_{1},t_{2}).$ (16)
Let $w$ stand for the corresponding vector field:
$w(t,z)=\Big(\frac{\partial H}{\partial p_{1}},\ldots,\frac{\partial H}{\partial p_{m}},-\frac{\partial H}{\partial x^{1}},\ldots,-\frac{\partial H}{\partial x^{m}}\Big).$ (17)
Observe that $w$ is uniquely determined from the equation
$i_{w}\beta=-d_{z}H.$ (18)
Therefore this equation can be taken as an invariant definition of a Hamiltonian vector field.
Together with system (16) it is convenient to study its autonomous version in the extended phase space
$\frac{dt}{d\tau}=1,\quad\frac{dx^{i}}{d\tau}=\frac{\partial H}{\partial p_{i}},\quad\frac{dp_{i}}{d\tau}=-\frac{\partial H}{\partial x^{i}}.$ (19)
The corresponding vector field is
$\tilde{w}(t,z)=\Big(1,\frac{\partial H}{\partial p_{1}},\ldots,\frac{\partial H}{\partial p_{m}},-\frac{\partial H}{\partial x^{1}},\ldots,-\frac{\partial H}{\partial x^{m}}\Big).$
We use $\tilde{N}$ to denote a phase space of system (19). The manifold $\tilde{N}$ is an extended phase space of (16):
$(t,z)\in\tilde{N}=I\times N.$
Let $g^{\tau}:\tilde{N}\to\tilde{N}$ stand for the flow of (19).
Let $G_{t_{0}}^{t}:N\to N$ denote the shift along the trajectories of (16):
$\frac{dG^{t}_{t_{0}}(z)}{dt}=w(t,G^{t}_{t_{0}}(z)),\quad G^{t_{0}}_{t_{0}}(z)=z.$
Recall that
$g^{\tau}\big((t_{0},z)\big)=\big(t_{0}+\tau,G^{t_{0}+\tau}_{t_{0}}(z)\big).$
Introduce a differential form on $\tilde{N}$
$\alpha=p_{i}dx^{i}-Hdt.$
This form is referred to as the Poincare-Cartan relative integral invariant. We justify this name below.
Theorem 11.
The following formula holds $i_{\tilde{w}}d\alpha=0$.
Conversely if a vector field $u(t,z)$ satisfies the equality $i_{u}d\alpha=0$ then $u=\lambda(t,z)\tilde{w}$.
Proof
Present $d\alpha$ as follows
$d\alpha=dp_{i}\wedge dx^{i}-dH\wedge dt=\pi_{i}\wedge\varkappa^{i},$
where
$\pi_{i}=dp_{i}+\frac{\partial H}{\partial x^{i}}dt,\quad\varkappa^{i}=dx^{i}-\frac{\partial H}{\partial p_{i}}dt.$
A value of the form $d\alpha$ for vectors $\xi,\eta$ is calculated as
$d\alpha(\xi,\eta)=\pi_{i}(\xi)\varkappa^{i}(\eta)-\pi_{i}(\eta)\varkappa^{i}(\xi).$
Since
$\varkappa^{i}(\tilde{w})=\pi_{i}(\tilde{w})=0,$
we have
$i_{\tilde{w}}d\alpha=\pi_{i}(\tilde{w})\varkappa^{i}(\cdot)-\pi_{i}(\cdot)\varkappa^{i}(\tilde{w})=0.$
This proves the first part of the theorem.
Let us check the second one:
$i_{u}d\alpha=\pi_{i}(u)\varkappa^{i}(\cdot)-\pi_{i}(\cdot)\varkappa^{i}(u)=0.$
Note that 11-forms $\pi_{i},\varkappa^{i},\quad i=1,\ldots,m$ are linearly independent so that
$\varkappa^{i}(u)=\pi_{i}(u)=0.$
Then putting $u=(u_{t},u_{x},u_{p})$ we find
$u_{p}+\frac{\partial H}{\partial x}u_{t}=0,\quad u_{x}-\frac{\partial H}{\partial p}u_{t}=0.$
The theorem is proved.
Theorem 12.
The form $\alpha$ is a relative integral invariant of system (19):
$L_{\tilde{w}}\alpha=d\mathscr{F},\quad\mathscr{F}=p_{i}\frac{\partial H}{\partial p_{i}}-H.$ (20)
Proof
We can check this by direct calculation from formula (4). But it is convenient to use the Cartan formula and theorem 11:
$L_{\tilde{w}}\alpha=di_{\tilde{w}}\alpha+i_{\tilde{w}}d\alpha,\quad di_{\tilde{w}}\alpha=d\mathscr{F}.$
The theorem is proved.
Observe that if the Hamiltonian $H$ admits the Legendre transform in the impulses then $\mathscr{F}$ is a Lagrangian function.
Theorem 13.
Let $\gamma\subset\tilde{N}$ be a closed curve. Then the integral
$\int_{g^{\tau}(\gamma)}\alpha$
does not depend on $\tau$.
This fact is a direct consequence from theorem 3.
Remark 2 implies that the form $d\alpha$ is an integral invariant: $L_{\tilde{w}}d\alpha=0.$
Theorem 14.
Let $\Sigma\subset\tilde{N}$ be an admissible two-dimensional manifold. Then the integral
$\int_{g^{\tau}(\Sigma)}d\alpha$
does not depend on $\tau.$
This fact is a direct consequence from theorem 2.
Theorem 15.
Let $\Sigma\subset N$ be an admissible two-dimensional manifold. Then the integral
$\int_{G^{t}_{t_{0}}(\Sigma)}\beta$
does not depend on $t,t_{0}$.
Proof
This fact is a direct consequence from theorem 6. Indeed, since $d_{z}\beta=0$ and by virtue of (18) we obtain
$\frac{\partial\beta}{\partial t}+L_{w}\beta=d_{z}i_{w}\beta+i_{w}d_{z}\beta=0.$
The theorem is proved.
Corollary 2.
From theorem 15 it follows that $G^{t}_{t_{0}}$ preserves the form $\beta:$
$(G^{t}_{t_{0}})_{*}\beta=\beta.$ (21)
In other words the shift along the trajectories of a Hamiltonian system is a symplectic map.
Moreover formula (21) implies that this shift preserves a volume in the phase space:
$(G^{t}_{t_{0}})_{*}\varpi=\varpi,\quad\varpi=\underbrace{\beta\wedge\ldots\wedge\beta}_{m\,\,\mbox{times}}.$
Theorem 16.
If $\gamma\subset N$ is a closed curve then the integral
$\int_{G^{t}_{t_{0}}(\gamma)}\zeta,\quad\zeta=p_{i}dx^{i}$
does not depend on $t,t_{0}$.
This fact is a direct consequence from theorem 6. Indeed:
$\frac{\partial\zeta}{\partial t}+L_{w}\zeta=d_{z}i_{w}\zeta+i_{w}d_{z}\zeta=d_{z}\Big(p_{i}\frac{\partial H}{\partial p_{i}}-H\Big).$
Problem 6.
Consider a Hamiltonian system with a Hamiltonian
$H(t,x,p),\quad x=(x^{1},\ldots,x^{m}),\quad p=(p_{1},\ldots,p_{m})\in\mathbb{R}^{m},\quad t\in\mathbb{R}.$
Assume this Hamiltonian to be homogeneous in the impulses:
$H(t,x,\lambda p)=\lambda H(t,x,p),\quad\forall\lambda>0.$
Recall that such functions satisfy the Euler equation:
$H=p_{i}\frac{\partial H}{\partial p_{i}}.$
Let $\gamma\subset N$ be an admissible one-dimensional manifold. Show that the integral
$\int_{G_{t_{0}}^{t}(\gamma)}p_{i}dx^{i}$
does not depend on $t,t_{0}$.
Problem 7.
Let the Hamiltonian $H$ be the same as in problem 6. Fix an initial condition $x(0)=\hat{x}$ and take an initial condition $p(0)=\hat{p}$ as a parameter.
[…]
$\frac{\partial x^{i}(t,\hat{p})}{\partial\hat{p}_{s}}p_{i}(t,\hat{p})=0,\quad s=1,\ldots,m.$ (22)
Hint: check the equality for $t=0$ and by direct calculation show that
$\frac{d}{dt}\Big(\frac{\partial x^{i}(t,\hat{p})}{\partial\hat{p}_{s}}p_{i}(t,\hat{p})\Big)=0.$
To conclude this observe that a form $p_{i}dt\wedge dx^{i}$ is a relative integral invariant for (19) and the form $dt\wedge dp_{i}\wedge dx^{i}$ is an integral invariant.
5.1. Reduction of the Order of a Hamiltonian System by Use of the Energy Integral
Let us point out a consequence from theorem 11.
Assume that the Hamiltonian $H$ does not depend on $t$. Then $H$ is a first integral. Assume that the energy level
$E_{h}=\{H(z)=h\}\subset\tilde{N}$
is not degenerate: $dH\mid_{E_{h}}\neq 0$.
Without loss of generality we can assume that in any small enough domain of the manifold $E_{h}$ the following inequality holds
$\frac{\partial H}{\partial p_{1}}\neq 0.$
We can always achieve this by the canonical permutations  [^1] Changes of variables $(x,p)\mapsto(X,P)$ of the type $X^{i}=-p_{i},\quad P_{i}=x^{i},\quad P_{s}=p_{s},\quad X^{s}=x^{s},\quad s\neq i$ or $X^{i}=x^{j},\quad X^{j}=x^{i},\quad P_{i}=p_{j},\quad P_{j}=p_{i}$ and compositions of such changes we call the canonical permutations. From the results of section 8 it follows that the canonical permutations preserve the Hamiltonian shape of equations..
Then the Implicit function theorem implies that the surface $E_{h}$ is described locally by the graph
$p_{1}=g(x^{1},\ldots,x^{m},p_{2},\ldots,p_{m},h),$ (23)
and $t,x^{1},\ldots,x^{m},p_{2},\ldots,p_{m}$ are local coordinates in $E_{h}$.
Therefore we obtain
$\alpha\mid_{E_{h}}=\sum_{k=2}^{m}p_{k}dx^{k}-\big(-g(x^{1},x^{2},\ldots,x^{m},p_{2},\ldots,p_{m},h)\big)dx^{1}-hdt.$
Theorem 11 and the equality $d(hdt)=0$ imply that the trajectories of (19) that belong to $E_{h}$ coincide with trajectories of the following system
$\frac{dt}{dT}=b(t,x^{1},\ldots,x^{m},p_{1},\ldots,p_{m}),\quad\frac{dx^{1}}{dT}=1,\quad\frac{dp_{i}}{dT}=\frac{\partial g}{\partial x^{i}},\quad\frac{dx^{i}}{dT}=-\frac{\partial g}{\partial p_{i}},\quad i=2,\ldots,m.$ (24)
From the second equation of system (24) we obtain
$x^{1}=T+T_{0},\quad T_{0}=\mathrm{const}.$
Thus system (24) separates in two sets of equations. The set of equations for variables
$x^{i},p_{i},\quad i=2,\ldots,m$
turns to be a Hamiltonian system with the Hamiltonian
$\mathscr{H}(T,x^{2},\ldots,x^{m},p_{2},\ldots,p_{m},h):=-g(T+T_{0},x^{2},\ldots,x^{m},p_{2},\ldots,p_{m},h).$
Since
$\dot{x}^{1}=\frac{\partial H}{\partial p_{1}}\neq 0$
we can parametrize a trajectory lying on $E_{h}$ by the parameter $x^{1}$. Then
$b=\frac{1}{\frac{\partial H}{\partial p_{1}}}.$
6. Characteristic Property of the Hamilton-Jacobi Equation
Theorem 17.
1) Assume that a function $S=S(t,x)$ satisfies the Hamilton-Jacobi equation:
$H\Big(t,x,\frac{\partial S}{\partial x}\Big)+\frac{\partial S}{\partial t}=0.$
Then the graph
$\Gamma=\Big\{p_{i}=\frac{\partial S}{\partial x^{i}}(t,x),\quad i=1,\ldots,m\Big\}\subset\tilde{N}$
is an $m+1$-dimensional invariant manifold of system (19).
2) Conversely, let a graph
$\Big\{p=\frac{\partial S}{\partial x}\Big\}\subset\tilde{N},\quad S=S(t,x)$
be an invariant surface of system (19). Then there exists a function $\psi=\psi(t)$ such that the function $S$ satisfies the equation
$H\Big(t,x,\frac{\partial S}{\partial x}\Big)+\frac{\partial S}{\partial t}=\psi(t).$
(Then a function $\tilde{S}=S-\int\psi dt$ satisfies the Hamilton-Jacobi equation.)
3) if $S$ is a solution to the Hamilton-Jacobi equation then $\alpha\mid_{\Gamma}=dS$.
Proof
The last assertion is trivial. Let us prove item 1).
Observe that $t,x$ are the local coordinates on the manifold $\Gamma$.
Let $x(t)$ be a solution to the system
$\dot{x}^{i}=\frac{\partial H}{\partial p_{i}}\Big(t,x,\frac{\partial S}{\partial x}(t,x)\Big).$
Show that
$x(t),\quad p(t)=\frac{\partial S}{\partial x}(t,x(t))$
is a solution to the Hamilton equations.
Indeed, introduce a function
$F(t,x)=H\Big(t,x,\frac{\partial S}{\partial x}\Big).$
Then we have
$\frac{\partial F}{\partial x^{i}}=\frac{\partial H}{\partial x^{i}}+\frac{\partial H}{\partial p_{s}}\frac{\partial^{2}S}{\partial x^{i}\partial x^{s}}$ (25)
and calculate
$\dot{p}_{k}$ $=\frac{\partial^{2}S}{\partial t\partial x^{k}}+\frac{\partial H}{\partial p_{s}}\frac{\partial^{2}S}{\partial x^{k}\partial x^{s}}=\frac{\partial^{2}S}{\partial t\partial x^{k}}+\frac{\partial F}{\partial x^{k}}-\frac{\partial H}{\partial x^{k}}$
$=\frac{\partial}{\partial x^{k}}\Big(\frac{\partial S}{\partial t}+F\Big)-\frac{\partial H}{\partial x^{k}}=-\frac{\partial H}{\partial x^{k}}.$
To prove 2) take a solution $x(t),p(t)$ to the Hamilton equations and differentiate the equality
$p_{i}(t)=\frac{\partial S}{\partial x^{i}}(t,x(t))$
in $t$:
$\frac{\partial H}{\partial x^{i}}=\frac{\partial^{2}S}{\partial x^{i}\partial x^{r}}\frac{\partial H}{\partial p_{r}}+\frac{\partial^{2}S}{\partial x^{i}\partial t}.$
Using formula (25) we yield
$\frac{\partial}{\partial x^{i}}\Big(\frac{\partial S}{\partial t}+F\Big)=0.$
The theorem is proved.
Item 3) implies two consequences.
Firstly we have $d\alpha\mid_{\Gamma}=0.$
Secondly if $(\tau,x(\tau),p(\tau))$ is a trajectory of (19) that belongs to the manifold $\Gamma$ then the following equality holds
$S(t,x(t))-S(t_{0},x(t_{0}))$ $=\int_{t_{0}}^{t}\big(p_{i}(\tau)\dot{x}^{i}(\tau)-H(\tau,(x(\tau),p(\tau))\big)d\tau$
$=\int_{t_{0}}^{t}L(\tau,x(\tau),\dot{x}(\tau))d\tau,$ (26)
where $L$ is a Lagrangian function.
The second equality holds only if the Legendre transform of $H$ is correctly defined. Indeed,
$p_{i}(\tau)\dot{x}^{i}(\tau)-H(\tau,(x(\tau),p(\tau))=p_{i}(\tau)\frac{\partial H}{\partial x^{i}}-H(\tau,(x(\tau),p(\tau)).$
For systems of Classical mechanics there are no problems with the Legendre transform.
Note also that if the Hamiltonian $H$ is a homogeneous function in $p$ like in Problem 6 then the first line of equality (26) takes the form
$S(t,x(t))=S(t_{0},x(t_{0})).$
7. Eikonal Equation and Gauss’s Lemma
7.1. Eikonal Equation
Let $M$ be a Riemannian manifold with local coordinates
$x=(x^{1},\ldots,x^{m})$
and let $g_{ij}(x)$ stand for a metric tensor.
Let $x(t)$ denote a solution to the Lagrange equations with a Lagrangian
$L=\frac{1}{2}g_{ij}(x)\dot{x}^{i}\dot{x}^{j},\quad H=\frac{1}{2}g^{ij}(x)p_{i}p_{j},$ (27)
and put
$x(0)=\hat{x},\quad p_{i}=\frac{\partial L}{\partial\dot{x}^{i}}=g_{ij}(x)\dot{x}^{j}.$
We assume that the solution $x(t)$ has energy $1/2$:
$|\dot{x}(t)|^{2}=g_{ij}(x(t))\dot{x}^{i}(t)\dot{x}^{j}(t)=1.$ (28)
Then the equation $x=x(t)$ defines a geodesic and $t$ is an arc-length parameter.
Theorem 18.
Let a function $f:M\to\mathbb{R}$ satisfy the eikonal equation
$|\nabla f|^{2}=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial f}{\partial x^{j}}=1,\quad\nabla f=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}.$ (29)
Then if a geodesic $x(t)$ starts from a level surface
$\Psi_{\hat{x}}=\{x\in M\mid f(x)=f(\hat{x})\},\quad\hat{x}\in\Psi_{\hat{x}}$
and this geodesic is perpendicular to this surface:
$\frac{\partial f}{\partial x^{i}}(\hat{x})=g_{ij}(\hat{x})\dot{x}^{j}(0),$ (30)
then
1) it is perpendicular to each surface $\Psi_{x(t)}$ it meets:
$\frac{\partial f}{\partial x^{i}}(x(t))=g_{ij}(x(t))\dot{x}^{j}(t);$
and
2) $f(x(t))-f(\hat{x})=t.$
Proof of Theorem 18
A function S(t,x)=f(x)t/2S(t,x)=f(x)-t/2 satisfies the Hamilton-Jacobi equation with Hamiltonian (27). Moreover due to condition (30) and theorem 17 the geodesic $x(t)$ belongs to the surface $\Gamma$:
$p_{i}(t)=g_{ij}(x(t))\dot{x}^{j}(t)=\frac{\partial S}{\partial x^{i}}(x(t))=\frac{\partial f}{\partial x^{i}}(x(t)).$
This means that the vector $\dot{x}(t)$ is perpendicular to the level surface $\Psi_{x(t)}$.
By virtue of formula (28) we have
$\int_{0}^{t}Ld\tau=\frac{t}{2},$
consequently from formula (26) we find
$S(t,x(t))-S(0,\hat{x})=\frac{t}{2}.$
This equality implies item 2) of the theorem.
Theorem 18 is proved.
Theorem 19.
Let $\Sigma\subset M$ be a hypersurface. Then locally in some neighborhood of this surface there exists a smooth function $f$ that solves equation (29) and satisfies the boundary conditions
$f\mid_{\Sigma}=\hat{f}=\mathrm{const}.$ (31)
Remark 5.
Cauchy problem (29), (31) does not possess uniqueness but this problem has only two solutions. This follows from item 2) of theorem 18.
Proof of Theorem 19
Choose an arbitrary point $\hat{x}\in\Sigma$ and introduce in its neighborhood $U$ local coordinates $x$ such that $U\cap\Sigma=\{x^{1}=0\}.$ Equation
$g^{ij}(x)s_{i}s_{j}=1$ (32)
has a solution $s_{1}=\hat{s}_{1}=\frac{1}{\sqrt{g^{11}(\hat{x})}},\quad s_{2}=\ldots=s_{m}=0,\quad x=\hat{x}$, and
$\frac{\partial}{\partial s_{1}}g^{ij}(x)s_{i}s_{j}\Big|_{s_{2}=\ldots,=s_{m}=0,\,\,x=\hat{x},\,\,s_{1}=\hat{s}_{1}}=2\sqrt{g^{11}(\hat{x})}\neq 0.$
From the Implicit function theorem we conclude that locally equation (32) has a solution:
$s_{1}=\varphi(x,s_{2},\ldots,s_{m}).$
Therefore equation (29) is locally equivalent to the following one
$\frac{\partial f}{\partial x^{1}}=\varphi\Big(x,\frac{\partial f}{\partial x^{2}},\ldots,\frac{\partial f}{\partial x^{m}}\Big).$
Condition (31) takes the form
$f\mid_{x^{1}=0}=\hat{f}$.
This Cauchy problem is solved by the method of characteristics (see Section 11) for $|x^{1}|$ sufficiently small.
The theorem is proved.
Problem 8.
Let a function $f:M\to\mathbb{R}$ satisfy equation (29). Show that all the solutions to ODE
$\dot{x}=\nabla f$ (33)
are the geodesics.
Solution to Problem 8
The solution repeats the argument above almost literally.
We employ the same solution to the Hamilton-Jacobi equation: S(t,x)=t/2+f(x)S(t,x)=-t/2+f(x).
[…]
Let $(x,p)(t)$ be a solution to the Hamilton equation with the initial conditions $\hat{x},\hat{p}$. This particularly implies that $x(t)$ is a solution to the Lagrange equations and thus $x(t)$ is a geodesic.
By theorem 17 we have
$p_{i}(t)=g_{ij}(x(t))\dot{x}^{j}(t)=\frac{\partial S}{\partial x^{i}}=\frac{\partial f}{\partial x^{i}}.$
So that $x(t)$ is a solution to (33).
But there is a shorter way to solve problem 8. Any solution to equation (33) provides a minimum of the Action functional :
$\int_{t_{1}}^{t_{2}}\frac{1}{2}|\dot{x}-\nabla f|^{2}dt.$
Therefore any solution to equation (33) is a solution to the Lagrange equations with the Lagrangian
$L=\frac{1}{2}|\dot{x}-\nabla f|^{2}.$
This Lagrangian coincides with
$\frac{1}{2}|\dot{x}|^{2}$
up to a total derivative.
Problem 9.
Let
$f(x,a),\quad a=(a^{1},\ldots,a^{\ell})\in\mathbb{R}^{\ell}$
be an $\ell$-parametric family of solutions to equation (29).
Assume that the equation
$\frac{\partial f}{\partial a}(x,a)=0$
has a solution: $a=\theta(x)$.
Show that the envelop
$\tilde{f}(x):=f(x,\theta(x))$
is a solution to equation (29) as well [5].
Show that for all possible $a$ the surface $\{\tilde{f}(x)=c\}$ is tangent to each surface of the family $\{f(x,a)=c\}$.
7.2. Gauss’s Lemma
Consider the result of problem 7 from Riemannian geometry view point.
A Hamiltonian
$\mathcal{H}=\sqrt{g^{ij}(x)p_{i}p_{j}}$
is a homogeneous function in the impulses. The function $\mathcal{H}$ is a first integral of the system. Let $x(t),p(t)$ be a solution with $x(0)=\hat{x}$ fixed and such that
$\mathcal{H}=1.$ (34)
We regard the initial value $p(0)=\hat{p}$ as a parameter.
It is easy to see that such solutions satisfy the system with Hamiltonian (27) as well. Condition (34) is equivalent to (28). Particularly $t$ is an arc-length parameter and the solutions $x(t)$ of the described above type are the geodesics on $M$ and these geodesics run from the point $\hat{x}$.
Introduce a sphere
$C=\{v=(v^{1},\ldots,v^{m})\in T_{\hat{x}}M\mid|v|^{2}=g_{ij}(\hat{x})v^{i}v^{j}=1\}.$
A vector $v\in C$ is put in correspondence with a geodesic $x_{v}(t),\quad x_{v}(0)=\hat{x},\quad x_{v}(0)=v$.
Take $\tau>0$ so small that different geodesics from the family $\{x_{v}\}_{v\in C}$ do not intersect with each other when $t\in(0,\tau]$ and a set $\Sigma=\{x_{v}(\tau)\in M\mid v\in C\}$ forms a smooth manifold of dimension m1m-1.
Theorem 20.
A vector $\dot{x}_{v}(\tau)\in T_{x_{v}(\tau)}M$ is perpendicular to the manifold $\Sigma$.
Indeed, take an arbitrary smooth curve
$x=x_{v(\xi)}(\tau),\quad|v(\xi)|=1$
on $\Sigma$. Here $\xi\in\mathbb{R}$ is a parameter on the curve.
A vector
$\frac{dx^{i}_{v(\xi)}(\tau)}{d\xi}=\frac{\partial x^{i}_{v}}{\partial v^{r}}\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}\frac{\partial\hat{p}_{j}}{\partial v^{r}}\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}g_{jr}(\hat{x})\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}\frac{d\hat{p}_{j}}{d\xi}.$
is a tangent vector to $\Sigma$.
Equality (22) is read as
$\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}g_{il}(x_{v}(\tau))\dot{x}^{l}_{v}(\tau)=0.$
So that
$g_{il}(x_{v}(\tau))\dot{x}^{l}_{v}(\tau)\frac{d}{d\xi}x^{i}_{v}(\tau)=0.$
The theorem is proved.
8. Canonical Transformations. Generating Functions
8.1. Canonical Transformations
Let $f=f(t,x,p)$ be a function defined on the extended phase space $\tilde{N}$ of system (16). Recall the notation:
$d_{z}f=\frac{\partial f}{\partial x^{i}}dx^{i}+\frac{\partial f}{\partial p_{i}}dp_{i},\quad df=d_{z}f+\frac{\partial f}{\partial t}dt.$
Definition 3 ([1]).
A transformation222we use the word ”transformation” as a synonym of the phrase ”change of variables” of the manifold $\tilde{N}$
$(t,x,p)\mapsto(t,X,P),\quad X=X(t,x,p),\quad P=P(t,x,p)$
is said to be canonical if
$d_{z}P_{i}\wedge d_{z}X^{i}=dp_{i}\wedge dx^{i}.$ (35)
Here and below we use $t,x,p$ as independent variables so that
$d_{z}p_{i}=dp_{i},\quad d_{z}x^{i}=dx^{i},\quad d_{z}t=0.$
Loosely speaking the canonical transformation is a family of symplectic maps parametrized with the parameter $t$. For example in accordance with Corollary 2 the mapping $(t,x,p)\mapsto(t,G^{t}_{t_{0}}(x,p))$ is a canonical transformation.
Formula (35) is locally equivalent to the following one
$p_{i}dx^{i}-P_{i}d_{z}X^{i}=d_{z}S(t,x,p)$ (36)
where $S$ is a function. Indeed:
$d_{z}\big(p_{i}dx^{i}-P_{i}d_{z}X^{i}\big)=0.$
Equality (36) is rewritten in the shape
$P_{i}dX^{i}+P_{i}\frac{\partial X^{i}}{\partial t}dt+p_{i}dx^{i}=dS-\frac{\partial S}{\partial t}dt.$ (37)
Theorem 21.
In the coordinates $(X,P)$ the Hamiltonian system (16) preserves its Hamiltonian form:
$\dot{P}_{i}=-\frac{\partial K}{\partial X^{i}},\quad\dot{X}^{i}=\frac{\partial K}{\partial P_{i}},$
where
$K(t,X,P)=\Big(P_{i}\frac{\partial X^{i}}{\partial t}+\frac{\partial S}{\partial t}+H\Big)\Big|_{(x,p)\mapsto(X,P)}.$
Indeed, formula (37) implies
$\alpha=p_{i}dx^{i}-Hdt=P_{i}dX^{i}-Kdt+dS,$ (38)
and the assertion of theorem 21 follows from theorem 11.
Corollary 3.
If a canonical transformation does not depend on time:
$P=P(x,p),\quad X=X(x,p),$
then the Hamiltonian $H$ is transformed as a function on the manifold $N$:
$K(t,X,P)=H(t,x(X,P),p(X,P)).$
8.2. Generating Functions
Definition 4.
A canonical transformation $(t,x,p)\mapsto(t,X,P)$ is said to be free if
$\det\Big(\frac{\partial X^{i}}{\partial p_{j}}\Big)\neq 0.$ (39)
In this case by the Implicit function theorem the variables $(t,x,X)$ are local coordinates in $\tilde{N}$. In particular we can write $S=S_{1}(t,x,X).$
Formula (38) takes the form
$p_{i}dx^{i}-Hdt=P_{i}dX^{i}-Kdt+\frac{\partial S_{1}}{\partial t}dt+\frac{\partial S_{1}}{\partial x^{i}}dx^{i}+\frac{\partial S_{1}}{\partial X^{i}}dX^{i}$ (40)
and
$p_{i}$ $=\frac{\partial S_{1}}{\partial x^{i}},\quad P_{i}=-\frac{\partial S_{1}}{\partial X^{i}};$ (41)
$K$ $=H+\frac{\partial S_{1}}{\partial t}.$ (42)
Inequality (39) is written as follows
$\det\Big(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\Big)\neq 0.$ (43)
Conversely if inequality (43) is fulfilled then formulas (41) define a canonical transformation.
Indeed, let us check the nondegenerateness of the change $(x,p)\mapsto(X,P)$.
Present this change as a composition
$(x,p)\mapsto(x^{\prime},X^{\prime})\mapsto(X,P).$
The first change is the inverse for the change
$x=x^{\prime},\quad p=\frac{\partial S_{1}}{\partial x}(t,x^{\prime},X^{\prime}).$
The Jacobian of this change is
$\det\Big(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\Big)^{-1}.$
The second change is given by the formulas
$X=X^{\prime},\quad P=-\frac{\partial S_{1}}{\partial X}(t,x^{\prime},X^{\prime}).$
Its Jacobian is
$\det\Big(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\Big).$
Therefore if the function $S_{1}$ satisfies (43) then it defines a free change of variables which is restored by the formulas (41). The function $S_{1}$ is referred to as a generating function.
In the new variables $(X,P)$ the Hamiltonian $K$ is given by the formula (42).
Remark 6.
Assume that we know a generating function $S_{1}(t,x,X)$ that satisfies the Hamilton-Jacobi equation for each admissible $X$:
$H\Big(t,x,\frac{\partial S_{1}}{\partial x}\Big)+\frac{\partial S_{1}}{\partial t}=0.$
Then we know the change $(t,x,p)\mapsto(t,X,P)$ and in the new variables $X,P$ the Hamiltonian $K=0$ so that the system is trivially integrated.
Consider a composition of canonical transformations
$(x,p)\mapsto(\tilde{X},\tilde{P})\mapsto(X,P).$
The first transformation is defined by a generating function $S_{1}$:
[…]
and the second one is a canonical permutation:
$\tilde{X}=P,\quad\tilde{P}=-X.$
Thus the change $(x,p)\mapsto(X,P)$ is expressed by means of a generating function
$S_{2}(t,x,P):=S_{1}(t,x,P).$
Formulas (44) imply:
$p_{i}=\frac{\partial S_{2}}{\partial x^{i}},\quad X^{i}=\frac{\partial S_{2}}{\partial P_{i}};$
condition (43) takes the form:
$\det\Big(\frac{\partial^{2}S_{2}}{\partial x^{i}\partial P_{j}}\Big)\neq 0;$
and condition (39) is written as follows:
$\det\Big(\frac{\partial P_{i}}{\partial p_{j}}\Big)\neq 0.$
By subjecting to a canonical permutation only several pairs of conjugate variables $\tilde{X}^{i},\tilde{P}_{i}$ and using the same argument we obtain $2^{m}$ different types of generating functions.
Remark 7.
The identity mapping
$P=p,\quad X=x$
is obtained with the help of $S_{2}$ generating function: $S_{2}=P_{i}x^{i}.$
Definition 5.
A function $S=S(t,x,b)$ that depends on $m$ parameters $b=(b_{1},\ldots,b_{m})$ is called the complete integral of the Hamilton-Jacobi equation
$H\Big(t,x,\frac{\partial S}{\partial x}\Big)+\frac{\partial S}{\partial t}=0$
provided it is a solution of the Hamilton-Jacobi equation for all $b$ and
$\det\Big(\frac{\partial^{2}S}{\partial x\partial b}\Big)\neq 0.$
Therefore if a complete integral is given then the system of Hamilton equations is integrated in the closed form. See Remark 6.
The vector $b$ can be interpreted as a set of new coordinates and we then obtain the function $S_{1}$; if we take $b$ as a set of new impulses then we have $S_{2}$; the vector $b$ can consist partly of new coordinates and partly of new impulses as well.
Today most of the known integrable Hamiltonian systems have been integrated due to a separation of variables procedure. For the autonomous case ($H=H(z)$) this means that in some suitable coordinates the Hamilton-Jacobi equation
$H\Big(x,\frac{\partial S}{\partial x}\Big)=K(b)$
admits a complete integral of the form
$S(x,b)=\sum_{k=1}^{m}S_{k}(x^{k},b).$
9. Hamiltonian Vector Field Straightening Theorem
Assume that system (16) is autonomous: $H=H(z).$
Theorem 22.
Assume that the Hamiltonian $H$ is nondegenerate at a point $\tilde{z}\in N:$
$dH(\tilde{z})\neq 0.$
Then in some neighborhood $U$ of the point $\tilde{z}$ there exist coordinates
$Z=(X,P)=(X^{1},\ldots,X^{m},P_{1},\ldots,P_{m}),\quad Z=Z(x,p)$
such that
1) $dp_{i}\wedge dx^{i}=dP_{i}\wedge dX^{i};$
2) in coordinates $Z$ the Hamiltonian has the form $H=X^{1}.$
Proof
Without loss of generality put $\tilde{z}=0$ and
$H(0)=0,\quad\frac{\partial H}{\partial p_{1}}(0)\neq 0.$ (45)
Then by the Implicit function theorem an equation
$H(x,p)=X^{1}$
has a solution
$p_{1}=\phi(x,p_{2},\ldots,p_{m},X^{1}),\quad H(x,\phi,p_{2},\ldots,p_{m})=X^{1}$
for $|X^{1}|,|x|,|p|$ sufficiently small. (Cf. (23).) From (45) we get
$\phi(0)=0,\quad\frac{\partial H}{\partial p_{1}}(0)\cdot\frac{\partial\phi}{\partial X^{1}}(0)=1.$ (46)
Consider the following Cauchy problem:
$\frac{\partial S}{\partial x^{1}}=\phi\Big(x,\frac{\partial S}{\partial x^{2}},\ldots,\frac{\partial S}{\partial x^{m}},X^{1}\Big),\quad S\mid_{x^{1}=0}=\sum_{k=2}^{m}x^{k}X^{k},\quad S=S(x,X).$ (47)
From Section 11 we know that this Cauchy problem has a solution $S=S(x,X)$ for |x|,|X||x|,\,|X| small enough.
Thus equation (47) is equivalent to the equation
$H\Big(x,\frac{\partial S}{\partial x}\Big)=X^{1}.$
To finish the proof it remains to check that $S(x,X)$ is a generating function of a transformation $(x,p)\mapsto(X,P);$
$p=\frac{\partial S}{\partial x},\quad P=-\frac{\partial S}{\partial X}.$
The values $x=0,\quad p=0$ correspond to the values $X=0,\quad P=0$.
By virtue of (47) we have
$\frac{\partial^{2}S}{\partial x\partial X}\Big|_{x=X=0}=\begin{bmatrix}\frac{\partial\phi}{\partial X^{1}}&\frac{\partial^{2}S}{\partial x^{1}\partial X^{2}}&\dots&\frac{\partial^{2}S}{\partial x^{1}\partial X^{m}}\\ 0&1&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&1\end{bmatrix},\quad\det\frac{\partial^{2}S}{\partial x\partial X}\Big|_{x=X=0}=\frac{\partial\phi}{\partial X^{1}}.$
By formula (46) it follows that
$\frac{\partial\phi}{\partial X^{1}}(0)\neq 0.$
The theorem is proved.
10. Poincare Cut on the Energy Level
In this section we present a local description of a mutual disposition of such geometric objects as a Hamiltonian vector field, an energy level and an symplectic structure.
Alternative proofs of the following two theorems are contained in [9].
We use
$E_{h}=\{H(z)=h\}\subset N$
to denote an energy level; $dH\mid_{E_{h}}\neq 0$.
Recall that the Hamiltonian vector field $w$ is defined by formula (17).
Let a hypersurface
$Y\subset E_{h},\quad\dim Y=2m-2$
be such that $w(z)\notin T_{z}Y$ for all $z\in Y$.
Introduce a restriction $\beta^{\prime}=\beta\mid_{Y}.$
Theorem 23.
The form $\beta^{\prime}$ is nondegenerate. So $(Y,\beta^{\prime})$ is a symplectic manifold.
Proof
Take a point $z\in Y$. In a vicinity of this point introduce canonical coordinates in accordance with theorem 22. In these coordinates the integral curves are
$P_{1}(t)=-t+P_{1}(0),\quad P_{j}(t)=P_{j}(0),\quad X(t)=X(0),\quad j=2,\ldots,m$
and $E_{h}=\{X^{1}=h\}.$
Thus $X^{2},\ldots,X^{m},P_{1},\ldots,P_{m}$ are local coordinates in $E_{h}$ and the manifold $Y$ can be presented as a graph $P_{1}=y(X^{2},\ldots,X^{m},P_{2},\ldots,P_{m}).$ This implies that $(X^{2},\ldots,X^{m},P_{2},\ldots,P_{m})$ are local coordinates in $Y$.
The proof is accomplished by the following computation:
$\beta\mid_{Y}=dP_{i}\wedge dX^{i}\mid_{Y}=\sum_{k=2}^{m}dP_{k}\wedge dX^{k}.$
The theorem is proved.
Let $G^{t}:N\to N$ stand for a phase flow of (16). We know that $G^{t}_{*}\beta=\beta.$
Assume that manifolds
$Y_{1},Y_{2}\subset E_{h},\quad Y_{1}\cap Y_{2}=\emptyset$
have the same properties as the manifold $Y$ has.
As we proved above the manifolds
$(Y_{i},\beta_{i}),\quad\beta_{i}=\beta\mid_{Y_{i}}$
are symplectic.
Suppose that if a trajectory starts from a point $z\in Y_{1}$ then it reaches the surface $Y_{2}$ in time $t=\tau(z)>0$ so that
$Q(z):=G^{\tau(z)}(z)\in Y_{2}.$
Theorem 24.
The mapping $Q:Y_{1}\to Y_{2}$ is a symplectic mapping: $Q_{*}\beta_{2}=\beta_{1}$.
The proof of this theorem is trivial if the manifolds $Y_{1},Y_{2}$ are so close to each other that the arc $\{G^{t}(z)\mid t\in[0,\tau(z)]\}$ is contained in a domain with local coordinates $(X,P)$ from the proof of theorem 23.
If it is not the case we can take several manifolds
$\tilde{Y}_{1},\ldots,\tilde{Y}_{n}\subset E_{h}$ (48)
between the manifolds $Y_{1},Y_{2}$ and put
$\tilde{Y}_{0}:=Y_{1},\quad\tilde{Y}_{n+1}:=Y_{2}.$
The manifolds (48) have the same properties as the manifold $Y$ has and they are chosen such that the neighboring manifolds of the collection
$\tilde{Y}_{k},\quad k=0,1,\ldots,n+1$
are so close to each other that a piece of trajectory connecting the points
$z\in\tilde{Y}_{k},\quad G^{\tau_{k}(z)}(z)\in\tilde{Y}_{k+1}$
is contained in a domain with the coordinates straightening the vector field $w$.
Then $Q$ is a symplectic mapping as a chain of the symplectic mappings
$Q=Q_{n}\circ\ldots\circ Q_{0}$
where
$Q_{j}(z)=G^{\tau_{j}(z)}(z),\quad Q_{j}:\tilde{Y}_{j}\to\tilde{Y}_{j+1},\quad j=0,\ldots,n.$
Each manifold $\tilde{Y}_{k}$ has a symplectic form $\beta\mid_{\tilde{Y}_{k}}.$
By the same reason the result of theorem 24 remains valid if $Y_{1}=Y_{2}$.
11. The Hamilton-Jacobi Equation in General Case: the Method of Characteristics
In the theory of PDE the equation
$u_{t}+f\Big(t,x,u,\frac{\partial u}{\partial x}\Big)=0$ (49)
is also called the Hamilton-Jacobi equation.
Here $f=f(t,x,\xi,p)$ is a scalar function of the variables
$t,\xi\in\mathbb{R},\quad p=(p_{1},\ldots,p_{m}),\,x=(x^{1},\ldots,x^{m})\in\mathbb{R}^{m}.$
Consider a system of ODE:
$\dot{\xi}$ $=p_{i}\frac{\partial f}{\partial p_{i}}-f,$ (50)
$\dot{p}_{i}$ $=-\frac{\partial f}{\partial x^{i}}-\frac{\partial f}{\partial\xi}p_{i},$
$\dot{x}^{i}$ $=\frac{\partial f}{\partial p_{i}}.$
An extended phase space $V$ of this system is a space of variables
$V=\{(t,x,\xi,p)\in\mathbb{R}^{2m+2}\}.$
If the function $f$ does not depend on $\xi$ then the equations for $x,p$ are separated in (50) and form a system of Hamiltonian equations while equation (49) becomes a Hamilton-Jacobi equation in the sense regarded above.
System (50) is called the characteristic system and its solutions are called the characteristics.
The proof of the following theorem is the same as the proof of theorem 17.
Theorem 25.
Let $u=u(t,x)$ be a solution to (49). Then $m+1$ dimensional manifold (graph)
$G=\Big\{\xi=u(t,x),\quad p_{i}=\frac{\partial u}{\partial x^{i}}(t,x),\qquad i=1,\ldots,m\Big\}\subset V$
is an invariant manifold of system (50) and the following formula holds
$(p_{i}dx^{i}-fdt-d\xi)\mid_{G}=0$
or equivalently
$(p_{i}dx^{i}-fdt)\mid_{G}=du.$
This theorem shows a way to solve a Cauchy problem for equation (49). Indeed, provide this equation with initial data
$u\mid_{t=0}=\hat{u}(x).$ (51)
Define initial conditions for system (50):
[…]
Then a solution $u(t,x)$ to Cauchy problem (49), (51) is found from the equations
[…]
By the Implicit function theorem the last equation can be solved with respect to $\hat{x}$ for $|t|$ small enough.
Such a method of Cauchy problem (49), (51) solving is referred to as the method of characteristics.