### v9

# Lecture Notes in Integral Invariants and Hamiltonian Systems

Abstract.
In this methodological review, we discuss the fundamental concepts of the theory of integral invariants. This theory originated with Poincare and Cartan [8, 2] and was further developed by Kozlov [6]. We demonstrate how the core ideas of this theory link diverse fields of mathematical physics, such as Hamiltonian dynamics, optics, and hydrodynamics. Particular attention is paid to results that are rarely expounded in standard textbooks.
1. Invariant Differential Forms
Consider a smooth dynamical system
$\dot{x}=v(x),$ (1)
defined on the smooth manifold $M$ with local coordinates $x=(x^{1},\dots,x^{m})^{T}$.
Remark 1.
Throughout this text, all objects are assumed to be sufficiently smooth for the respective expressions to be well-defined, and all integrals are assumed to exist. Whenever Stokes’ theorem is applied, we assume that the necessary regularity and boundary conditions are satisfied.
We denote the phase flow of this system by $g^{t}:M\to M$, which satisfies:
$\frac{d}{dt}g^{t}(\hat{x})=v(g^{t}(\hat{x})),\quad g^{0}(\hat{x})=\hat{x}$.
Recall the group property $g^{t+s}=g^{t}\circ g^{s}$.
For small $t$, we employ the following standard expansion:
$g^{t}(x)=x+v(x)t+o(t),\quad t\to 0.$ (2)
Definition 1 ([7]).
Consider a differential $k$-form
$\omega=\sum_{i_{1}<\dots<i_{k}}\omega_{i_{1}\dots i_{k}}(x)dx^{i_{1}}\wedge\dots\wedge dx^{i_{k}}.$
The Lie derivative of $\omega$ along the vector field $v$, denoted by $L_{v}\omega$, is defined as the following differential operator:
$L_{v}\omega=\left.\frac{d}{dt}\right|_{t=0}g^{t}_{*}\omega=\left.\frac{d}{dt}\right|_{t=0}\sum_{i_{1}<\dots<i_{k}}\omega_{i_{1}\dots i_{k}}(g^{t}(x))d(g^{t}(x))^{i_{1}}\wedge\dots\wedge d(g^{t}(x))^{i_{k}}.$
Recall the Cartan homotopy formula:
$L_{v}\omega=di_{v}\omega+i_{v}d\omega.$
The operator $i_{v}\omega$, defined as
$i_{v}\omega=\sum_{i_{2}<\dots<i_{k}}\omega_{ii_{2}\dots i_{k}}v^{i}dx^{i_{2}}\wedge\dots\wedge dx^{i_{k}},$
is called the interior product (or contraction) of the form $\omega$ with the vector field $v$.This operator satisfies the following antiderivation property:
$i_{v}(\omega\wedge\nu)=(i_{v}\omega)\wedge\nu+(-1)^{k}\omega\wedge(i_{v}\nu),$ (3)
where $k$ is the degree of $\omega$.
Recall that the exterior derivative satisfies the same antiderivation property:
$d(\omega\wedge\nu)=(d\omega)\wedge\nu+(-1)^{k}\omega\wedge(d\nu).$
Problem 1.
By using (2) show that
$L_{v}\big(\omega_{i}dx^{i}\big)$ $=\Big(\frac{\partial\omega_{i}}{\partial x^{s}}v^{s}+\frac{\partial v^{s}}{\partial x^{i}}\omega_{s}\Big)dx^{i},$ (4)
$L_{v}\Big(\rho(x)dx^{1}\wedge\ldots\wedge dx^{m}\Big)$ $=\frac{\partial(\rho v^{i})}{\partial x^{i}}dx^{1}\wedge\ldots\wedge dx^{m},$
$L_{v}f$ $=v^{i}\frac{\partial f}{\partial x^{i}},\quad f:M\to\mathbb{R},$
$L_{v}d$ $=dL_{v}.$ (5)
Definition 2.
A $k$-form $\omega$ is called an integral invariant of system (1) if and only if $L_{v}\omega=0$.
A $k$-form $\omega$ is called a relative integral invariant of system (1) if there exists a (k1)(k-1)-form $\Omega$ such that $L_{v}\omega=d\Omega$.
Remark 2.
From formula (5), it follows that if the form $\omega$ is a relative integral invariant, then the form $d\omega$ is an (absolute) integral invariant.
Let $\Sigma\subset M$ denote a smooth $k$-dimensional submanifold of $M$.
Theorem 1.
The following identity holds:
$\left.\frac{d}{dt}\right|_{t=0}\int_{g^{t}(\Sigma)}\omega=\int_{\Sigma}L_{v}\omega.$ (6)
Proof
We prove this identity for the case where $\Sigma$ is covered by a single coordinate chart.
Let $u:D\to M$ denote the corresponding embedding, where $D\subset\mathbb{R}^{k}$ is an open domain such that $u(D)=\Sigma$.
Thus, by applying the change of variables formula, we obtain:
$\int_{g^{t}(\Sigma)}\omega=\int_{D}(g^{t}\circ u)_{*}\omega=\int_{D}(u_{*}\circ(g^{t})_{*})\omega,$
and
$\frac{d}{dt}\Big|_{t=0}\int_{D}(u_{*}\circ(g^{t})_{*})\omega=\int_{D}u_{*}\frac{d}{dt}\Big|_{t=0}(g^{t})_{*}\omega=\int_{D}u_{*}(L_{v}\omega).$
The theorem is proved.
Corollary 1.
Formula (6) implies
$\frac{d}{dt}\int_{g^{t}(\Sigma)}\omega=\frac{d}{ds}\Big|_{s=0}\int_{g^{t+s}(\Sigma)}\omega=\frac{d}{ds}\Big|_{s=0}\int_{g^{s}\big(g^{t}(\Sigma)\big)}\omega=\int_{g^{t}(\Sigma)}L_{v}\omega.$ (7)
Theorem 2.
If $\omega$ is an integral invariant, then for any admissible $k$-dimensional submanifold $\Sigma\subset M$ (see Remark 1), the integral
$\int_{g^{t}(\Sigma)}\omega$
is independent of $t$. The converse is also true.
Indeed, this follows directly from (7).
Theorem 3.
Assume that $\omega$ is a relative integral invariant. Then for any compact $k$-dimensional submanifold $\Sigma\subset M,\quad\partial\Sigma=\emptyset$ the following integral
$\int_{g^{t}(\Sigma)}\omega$
does not depend on $t$.
Indeed, from formula (7) and Stokes’ theorem, we have:
$\frac{d}{dt}\int_{g^{t}(\Sigma)}\omega=\int_{g^{t}(\Sigma)}L_{v}\omega=\int_{g^{t}(\Sigma)}d\Omega=0.$
This holds because $\partial g^{t}(\Sigma)=\emptyset$.
2. Invariant Differential Forms of Systems with First Integrals
2.1. The Case $\dim M=2$
If $m=2$ and $q$ is an invariant $2$-form of system (1):
$L_{v}q=0,$
then $i_{v}q$ is a closed form. Indeed, by Cartan’s formula:
$L_{v}q=di_{v}q+i_{v}dq.$
Having $dq=0$, we obtain $di_{v}q=0$.
Thus, locally there exists a function $f$ such that:
$df=i_{v}q.$
It follows that:
$L_{v}f=i_{v}df=i_{v}i_{v}q=0.$
Consequently, the function $f$ is a first integral.
2.2. Invariant Forms on Level Sets of First Integrals
Assume that system (1) has a first integral $F$ such that:
$L_{v}F=0,\quad dF\neq 0,$
and an invariant $m$-form $\omega:\quad L_{v}\omega=0$.
Without loss of generality, we can locally assume that:
$\frac{\partial F}{\partial x^{m}}\neq 0.$
Then $\omega$ can be written as:
$\omega=\rho(x)dx^{1}\wedge\dots\wedge dx^{m}=\rho(x)\left(\frac{\partial F}{\partial x^{m}}\right)^{-1}dx^{1}\wedge\dots\wedge dx^{m-1}\wedge dF.$
In invariant terms, this means that the form $\omega$ is represented as the following exterior product:
$\omega=\lambda\wedge dF,$
where $\lambda$ is an (m1)(m-1)-form defined up to an additive form $\gamma$ such that $\gamma\wedge dF=0$.
Proposition 1.
The restricted differential form
$\lambda|_{Z},\quad Z=\{F=\text{const}\},$
is an integral invariant for the restriction of system (1) to the manifold $Z$.
Indeed,
$L_{v}\omega=(L_{v}\lambda)\wedge dF+\lambda\wedge(L_{v}dF),\quad L_{v}(dF)=dL_{v}F=0.$
Therefore, we have $(L_{v}\lambda)\wedge dF=0$.
Let $e_{1},\dots,e_{m}$ be a basis in $T_{x}M$ such that $e_{1},\dots,e_{m-1}$ is a basis in $T_{x}Z$.
We consequently obtain:
$\big((L_{v}\lambda)\wedge dF\big)(e_{1},\dots,e_{m})=c(L_{v}\lambda)(e_{1},\dots,e_{m-1})\cdot dF(e_{m})=0,$
where $c$ is a non-zero constant related to the definition of the exterior product.
Since $dF(e_{m})\neq 0$, it follows that $(L_{v}\lambda)(e_{1},\dots,e_{m-1})=0$.
The proposition is proved.
Taking into account the results of Section 2.1, we see that if system (1) has m2m-2 independent first integrals and an invariant $m$-form:
$L_{v}\omega=0,\quad\omega=\rho(x)dx^{1}\wedge\dots\wedge dx^{m},\quad\rho(x)>0,$
then this system is integrable in closed form.
3. The Lie Derivative in the Nonautonomous case
Consider a vector field
$v=(v^{1},\ldots,v^{m})(t,x),\quad(t,x)\in\tilde{M}=(t_{1},t_{2})\times M.$
We use $G^{t}_{t_{0}}:M\to M$ to denote the shift along the trajectories of the system
$\dot{x}=v(t,x),\quad\frac{dG^{t}_{t_{0}}(x)}{dt}=v(t,G^{t}_{t_{0}}(x)),\quad G^{t_{0}}_{t_{0}}(x)=x,\quad G_{t_{1}}^{t}\circ G_{t_{0}}^{t_{1}}=G_{t_{0}}^{t}.$ (8)
Let $g^{\tau}:\tilde{M}\to\tilde{M}$ be the flow of the system
$\frac{dz}{d\tau}=\tilde{v}(z),\quad z=(t,x^{1},\ldots,x^{m})^{T},\quad\tilde{v}=(1,v^{1},\ldots,v^{m})^{T}.$ (9)
Here, $\tilde{M}$ denotes the extended phase space of (8).
The following equality holds:
$g^{\tau}(t_{0},x)=(t_{0}+\tau,G_{t_{0}}^{t_{0}+\tau}(x)).$
Consider a form
$\omega=\sum_{i_{1}<\ldots<i_{k}}\omega_{i_{1}\ldots i_{k}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}}$
and fix local coordinates. We can consider $\omega$ as a form on the manifold $\tilde{M}$ or as a form on the manifold $M$. In the latter case, $t$ is regarded as a parameter.
Introduce the following notations:
$\frac{\partial\omega}{\partial t}=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}(t,x)}{\partial t}dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}},\quad\frac{\partial v}{\partial t}=\left(\frac{\partial v^{1}}{\partial t},\ldots,\frac{\partial v^{m}}{\partial t}\right);$
and
$d$ $=d_{t}+d_{x},\quad d_{t}\omega=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}(t,x)}{\partial t}dt\wedge dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}};$
$d_{x}\omega$ $=\sum_{i_{1}<\ldots<i_{k}}\frac{\partial\omega_{i_{1}\ldots i_{k}}}{\partial x^{l}}dx^{l}\wedge dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k}},$
$L_{v}\omega$ $:=d_{x}i_{v}\omega+i_{v}d_{x}\omega.$
For instance, we have $d_{t}\omega=dt\wedge\frac{\partial\omega}{\partial t}$.
Theorem 4.
The following formula holds:
$L_{\tilde{v}}\omega=\frac{\partial\omega}{\partial t}+L_{v}\omega+(dt)\wedge i_{\frac{\partial v}{\partial t}}\omega.$
Proof of Theorem 4
Let us split the vector field into the following summands:
$\tilde{v}=e+v_{*},\quad e=(1,0,\ldots,0),\quad v_{*}=(0,v^{1},\ldots,v^{m}).$
Correspondingly, $L_{\tilde{v}}=L_{e}+L_{v_{*}}.$
Observe that
$L_{e}\omega=\frac{\partial\omega}{\partial t},\quad i_{v_{*}}\omega=i_{v}\omega.$
By the homotopy formula, we get:
$L_{v_{*}}\omega=i_{v_{*}}d_{x}\omega+i_{v_{*}}d_{t}\omega+d_{x}i_{v_{*}}\omega+d_{t}i_{v_{*}}\omega.$ (10)
Since
$i_{v_{*}}d_{x}\omega=i_{v}d_{x}\omega,\quad d_{x}i_{v_{*}}\omega=d_{x}i_{v}\omega,$
equality (10) takes the form:
$L_{v_{*}}\omega=L_{v}\omega+i_{v_{*}}d_{t}\omega+d_{t}i_{v_{*}}\omega.$
By formula (3), it follows that:
$i_{v_{*}}d_{t}\omega=i_{v_{*}}\left((dt)\wedge\frac{\partial\omega}{\partial t}\right)=-(dt)\wedge i_{v}\frac{\partial\omega}{\partial t},$
and
$d_{t}i_{v_{*}}\omega=(dt)\wedge\frac{\partial i_{v}\omega}{\partial t}=(dt)\wedge\left(i_{\frac{\partial v}{\partial t}}\omega+i_{v}\frac{\partial\omega}{\partial t}\right).$
This proves the theorem.
Theorem 5.
For any admissible $k$-dimensional submanifold $A\subset M$, one has:
$\frac{d}{dt}\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{G^{t}_{t_{0}}(A)}\left(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\right).$ (11)
Proof
Introduce the manifold $A_{t}=\{t\}\times A\subset\tilde{M}$, where $g^{\tau}(A_{t})=\{t+\tau\}\times G^{t+\tau}_{t}(A)\subset\tilde{M}$.
Using Theorem 4 and formula 7, we calculate:
$\frac{d}{d\tau}\int_{g^{\tau}(A_{t_{0}})}\omega$ $=\int_{g^{\tau}(A_{t_{0}})}L_{\tilde{v}}\omega=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}(L_{\tilde{v}}\omega)\big|_{t=t_{0}+\tau}$
$=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}\left(\frac{\partial\omega}{\partial t}+L_{v}\omega\right)\bigg|_{t=t_{0}+\tau}.$
To finish the proof, it remains to observe that
$\int_{g^{\tau}(A_{t_{0}})}\omega=\int_{G^{t_{0}+\tau}_{t_{0}}(A)}\omega\big|_{t=t_{0}+\tau}.$
The theorem is proved.
Remark 3.
After the change of variables $x=G^{t}_{t_{0}}(y)$, formula (11) takes the form:
$\frac{d}{dt}\int_{A}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\int_{A}\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\int_{A}(G^{t}_{t_{0}})_{*}\left(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\right).$
Since $A$ is an arbitrary $k$-dimensional manifold, we have:
$\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=(G^{t}_{t_{0}})_{*}\left(\frac{\partial\omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\omega(t,\cdot)\right).$ (12)
Theorem 6.
1) Let $A\subset M$ be an admissible $k$-dimensional submanifold and assume that
$\frac{\partial\omega}{\partial t}+L_{v}\omega=0.$
Then we have
$\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{A}\omega(t_{0},\cdot).$
2) Suppose that there exists a (k1)(k-1)-form
$\Omega=\sum_{i_{1}<\ldots<i_{k-1}}\Omega_{i_{1}\ldots i_{k-1}}(t,x)dx^{i_{1}}\wedge\ldots\wedge dx^{i_{k-1}}$
such that
$\frac{\partial\omega}{\partial t}+L_{v}\omega=d_{x}\Omega,$
and $A\subset M$ is a compact submanifold without boundary ($\partial A=\emptyset$). Then we have
$\int_{G^{t}_{t_{0}}(A)}\omega(t,\cdot)=\int_{A}\omega(t_{0},\cdot).$
Theorem 6 follows from Theorem 5 in the same manner as Theorems 3 and 2 follow from formula (7).
The theorem is proved.
3.1. A Remark on Solutions to a PDE
Formula (12) provides a method for solving the following Cauchy problem:
$\frac{\partial\omega}{\partial t}(t,x)+L_{v(t,x)}\omega(t,x)=0,\quad\omega\mid_{t=t_{0}}=\hat{\omega}$.
Indeed, from this formula it follows that
$\frac{d}{dt}(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=0.$
Since $(G^{t}_{t_{0}})_{*}\omega(t,\cdot)$ does not depend on $t$, we obtain
$(G^{t}_{t_{0}})_{*}\omega(t,\cdot)=\hat{\omega},\quad\text{whence}\quad\omega(t,x)=\big((G^{t}_{t_{0}})^{-1}\big)_{*}\hat{\omega}$.
4. Applications to Hydrodynamics
In this section, we set $M=\mathbb{R}^{3}$ and let $(x^{1},x^{2},x^{3})$ be the standard right-handed Euclidean frame. We also adopt the convention $d=d_{x}$.
Let $\mathbf{A}(t,x)=(A_{i})$ and $\mathbf{B}(t,x)=(B_{i})$ be vector fields in $\mathbb{R}^{3}$, and let $f(t,x)$ be a scalar function. There are standard correspondences:
$f\mapsto\omega^{3}_{f}=f\,dx^{1}\wedge dx^{2}\wedge dx^{3}$
and
$\mathbf{A}\mapsto\omega^{1}_{\mathbf{A}}=A_{i}dx^{i},\quad\mathbf{A}\mapsto\omega^{2}_{\mathbf{A}}=A_{1}dx^{2}\wedge dx^{3}+A_{2}dx^{3}\wedge dx^{1}+A_{3}dx^{1}\wedge dx^{2}.$
These correspondences can be expressed in terms of the Hodge star operator.
The following formulas are verified by direct calculation:
$df=\omega^{1}_{\mathrm{grad}\,f},\quad d\omega^{1}_{\mathbf{A}}=\omega^{2}_{\mathrm{curl}\,\mathbf{A}},\quad d\omega^{2}_{\mathbf{A}}=\omega^{3}_{\mathrm{div}\,\mathbf{A}}$
and
$i_{\mathbf{B}}\omega^{1}_{\mathbf{A}}=(\mathbf{A},\mathbf{B}),\quad i_{\mathbf{B}}\omega^{2}_{\mathbf{A}}=\omega^{1}_{\mathbf{A}\times\mathbf{B}},\quad i_{\mathbf{B}}\omega^{3}_{f}=f\omega^{2}_{\mathbf{B}}.$
These identities imply the following theorem.
Theorem 7.
The following formulas hold:
$\frac{\partial\omega^{1}_{\mathbf{A}}}{\partial t}+L_{v}\omega^{1}_{\mathbf{A}}$ $=\omega^{1}_{\frac{\partial\mathbf{A}}{\partial t}+(\mathrm{curl}\,\mathbf{A})\times v}+d(v,\mathbf{A}),$
$\frac{\partial\omega^{2}_{\mathbf{A}}}{\partial t}+L_{v}\omega^{2}_{\mathbf{A}}$ $=\omega^{2}_{\frac{\partial\mathbf{A}}{\partial t}+\mathrm{curl}\,(\mathbf{A}\times v)+v\,\mathrm{div}\,\mathbf{A}},$
$\frac{\partial\omega^{3}_{f}}{\partial t}+L_{v}\omega^{3}_{f}$ $=\omega^{3}_{\frac{\partial f}{\partial t}+\mathrm{div}\,(fv)}.$
Theorem 8.
Let $\gamma$, $\Sigma$, and $D\subset\mathbb{R}^{3}$ be a closed curve, a two-dimensional surface, and a domain, respectively.
1) If
$\frac{\partial\mathbf{A}}{\partial t}+(\mathrm{curl}\,\mathbf{A})\times v=\mathrm{grad}\,\psi,\quad\psi=\psi(t,x),$ (13)
then
$\int_{G_{t_{0}}^{t}(\gamma)}\omega^{1}_{\mathbf{A}(t,\cdot)}=\int_{\gamma}\omega^{1}_{\mathbf{A}(t_{0},\cdot)};$
2) If
$\frac{\partial\mathbf{A}}{\partial t}+\mathrm{curl}\,(\mathbf{A}\times v)+v\,\mathrm{div}\,\mathbf{A}=0,$ (14)
then
$\int_{G_{t_{0}}^{t}(\Sigma)}\omega^{2}_{\mathbf{A}(t,\cdot)}=\int_{\Sigma}\omega^{2}_{\mathbf{A}(t_{0},\cdot)};$
3) If
$\frac{\partial f}{\partial t}+\mathrm{div}\,(fv)=0,$ (15)
then
$\int_{G_{t_{0}}^{t}(D)}\omega^{3}_{f(t,\cdot)}=\int_{D}\omega^{3}_{f(t_{0},\cdot)}.$
This theorem is a consequence of Theorems 6 and 7.
Problem 2.
Prove the equality
$\mathrm{curl}\,(\mathbf{A}\times\mathbf{B})=[\mathbf{A},\mathbf{B}]+\mathbf{A}\,\mathrm{div}\,\mathbf{B}-\mathbf{B}\,\mathrm{div}\,\mathbf{A}.$
We use square brackets to denote the commutator of the vector fields:
$\mathbf{A},\mathbf{B}]=\frac{\partial\mathbf{A}}{\partial x}\mathbf{B}-\frac{\partial\mathbf{B}}{\partial x}\mathbf{A}.$
Hint: prove and use the formula
$L_{u}i_{v}-i_{v}L_{u}=i_{[u,v]}.$
Let $v(t,x)$ be the velocity field of an ideal fluid subject to potential forces, and assume its density depends on pressure only (a barotropic fluid).
Under these assumptions, equation (14) holds with $\mathbf{A}=\mathrm{curl}\,v$. The velocity field $v$ also satisfies equation (13) with $\mathbf{A}=v$ [3].
The famous Helmholtz and Kelvin theorems follow from Theorem 8. Furthermore, equation (15) corresponds to the continuity equation.
5. Darboux Theorem [4], [9]
In this section, we consider some nontrivial applications of the theory developed above.
Let $D\subset\mathbb{R}^{2m}=\{x=(x^{1},\dots,x^{2m})\}$ be an open neighbourhood of the origin.
Suppose a differential form
$\omega=\sum_{i<j}\omega_{ij}(x)dx^{i}\wedge dx^{j}$
defined in $D$ is closed and non-degenerate:
$d\omega=0,\quad\det(\omega_{ij}(x))\neq 0,\quad x\in D.$
Theorem 9.
In some open neighbourhood $D^{\prime}\subset D$ of the origin, there exist local coordinates $y=(y^{1},\dots,y^{2m})$ such that the form $\omega$ has constant coefficients:
$\omega=\sum_{i<j}\omega^{\prime}_{ij}dy^{i}\wedge dy^{j},\quad\omega^{\prime}_{ij}=\text{const}_{ij}.$
From linear algebra, it is known that there exists a linear transformation $y\mapsto(q^{1},\dots,q^{m},p_{1},\dots,p_{m})$ such that
$\omega=dp_{i}\wedge dq^{i}.$
Proof of Theorem 9
Introduce a constant form
$\omega_{1}=\omega|_{x=0}=\sum_{i<j}\omega_{ij}(0)dx^{i}\wedge dx^{j},\quad d\omega_{1}=0.$
Construct a family of differential forms as follows:
$\Omega(t,x)=t\omega_{1}+(1-t)\omega=\sum_{i<j}\Omega_{ij}(t,x)dx^{i}\wedge dx^{j},\quad t\in[0,1],\quad d_{x}\Omega=0.$
These forms are well-defined in $D$.
Each form $\Omega(t,0)=\omega|_{x=0}$ is non-degenerate, so all the forms $\{\Omega(t,x)\}$ are non-degenerate in some neighbourhood of the origin. Since $d(\omega_{1}-\omega)=0$, by the Poincare lemma, there exists a $1$-form $\alpha$ such that
$\omega_{1}-\omega=d\alpha,\quad\alpha=\alpha_{k}(x)dx^{k}.$
Let us choose $\alpha$ such that $\alpha|_{x=0}=0$.
Define a time-dependent vector field $v(t,x)$ as follows:
$i_{v(t,x)}\Omega(t,x)=-\alpha,\quad v^{i}(t,x)=-\Omega^{ij}(t,x)\alpha_{j}(x).$
Since $v(t,0)=0$, the corresponding family of shifts $x\mapsto G^{t}_{0}(x)$ is defined for small $|x|$ and for $t\in[0,1]$.
Using formula (12), we obtain
$\frac{d}{dt}(G^{t}_{0})_{*}\Omega(t,\cdot)=(G^{t}_{0})_{*}\left(\frac{\partial\Omega}{\partial t}(t,\cdot)+L_{v(t,\cdot)}\Omega(t,\cdot)\right).$
From Cartan’s formula, we get
$L_{v(t,\cdot)}\Omega(t,\cdot)=i_{v(t,x)}d_{x}\Omega(t,x)+d_{x}i_{v(t,x)}\Omega(t,x)=-d\alpha.$
On the other hand,
$\frac{\partial\Omega}{\partial t}=\omega_{1}-\omega=d\alpha.$
Thus, we have
$\frac{d}{dt}(G^{t}_{0})_{*}\Omega(t,\cdot)=0.$
Consequently, $(G^{1}_{0})_{*}\Omega(1,\cdot)=(G^{0}_{0})_{*}\Omega(0,\cdot)=\omega$, and the mapping $G^{1}_{0}$ takes the form $\omega_{1}$ to the form $\omega$.
The theorem is proved.
6. Integral Invariants of the Hamilton Equations
Recall that a manifold $N$ with $\dim N=2m$ is called a symplectic manifold if it is endowed with a non-degenerate closed 2-form, denoted by $\beta$. Non-degeneracy means that the equality $i_{a}\beta=0$ implies $a=0$.
Darboux’s theorem (see Section 5) states that any point of $N$ belongs to a chart with local coordinates
$z=(x^{1},\ldots,x^{m},p_{1},\ldots,p_{m})$
such that
$\beta=dp_{i}\wedge dx^{i}.$
These coordinates are called symplectic or canonical.
Consider the Hamiltonian system with the Hamiltonian function $H=H(t,x,p)$:
$\dot{x}^{i}=\frac{\partial H}{\partial p_{i}},\quad\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}},\quad t\in I=(t_{1},t_{2}).$ (16)
Let $w$ stand for the corresponding vector field:
$w(t,z)=\Big(\frac{\partial H}{\partial p_{1}},\ldots,\frac{\partial H}{\partial p_{m}},-\frac{\partial H}{\partial x^{1}},\ldots,-\frac{\partial H}{\partial x^{m}}\Big).$ (17)
Note that $w$ is uniquely determined by the equation
$i_{w}\beta=-d_{z}H.$ (18)
Therefore, this equation can be taken as an invariant definition of a Hamiltonian vector field.
Alongside system (16), it is convenient to study its autonomous version in the extended phase space:
$\frac{dt}{d\tau}=1,\quad\frac{dx^{i}}{d\tau}=\frac{\partial H}{\partial p_{i}},\quad\frac{dp_{i}}{d\tau}=-\frac{\partial H}{\partial x^{i}}.$ (19)
The corresponding vector field is
$\tilde{w}(t,z)=\Big(1,\frac{\partial H}{\partial p_{1}},\ldots,\frac{\partial H}{\partial p_{m}},-\frac{\partial H}{\partial x^{1}},\ldots,-\frac{\partial H}{\partial x^{m}}\Big).$
We use $\tilde{N}$ to denote the phase space of system (19). The manifold $\tilde{N}$ is the extended phase space of (16):
$(t,z)\in\tilde{N}=I\times N.$
Let $g^{\tau}:\tilde{N}\to\tilde{N}$ stand for the flow of (19).
Let $G_{t_{0}}^{t}:N\to N$ denote the shift along the trajectories of (16):
$\frac{dG^{t}_{t_{0}}(z)}{dt}=w(t,G^{t}_{t_{0}}(z)),\quad G^{t_{0}}_{t_{0}}(z)=z.$
Recall that
$g^{\tau}\big((t_{0},z)\big)=\big(t_{0}+\tau,G^{t_{0}+\tau}_{t_{0}}(z)\big).$
Introduce a differential form on $\tilde{N}$:
$\alpha=p_{i}dx^{i}-Hdt.$
This form is referred to as the Poincare–Cartan relative integral invariant. We justify this name below.
Theorem 10.
The following formula holds:
$i_{\tilde{w}}d\alpha=0.$
Conversely, if a vector field $u(t,z)$ satisfies the equality $i_{u}d\alpha=0$, then $u=\lambda(t,z)\tilde{w}$ for some scalar function $\lambda(t,z)$.
Proof
The 2-form $d\alpha$ can be represented as follows:
$d\alpha=dp_{i}\wedge dx^{i}-dH\wedge dt=\pi_{i}\wedge\varkappa^{i},$
where
$\pi_{i}=dp_{i}+\frac{\partial H}{\partial x^{i}}dt,\quad\varkappa^{i}=dx^{i}-\frac{\partial H}{\partial p_{i}}dt.$
The value of the form $d\alpha$ on the vectors $\xi,\eta$ is calculated as
$d\alpha(\xi,\eta)=\pi_{i}(\xi)\varkappa^{i}(\eta)-\pi_{i}(\eta)\varkappa^{i}(\xi).$
Since $\varkappa^{i}(\tilde{w})=0$ and $\pi_{i}(\tilde{w})=0$ by construction, we have
$i_{\tilde{w}}d\alpha=\pi_{i}(\tilde{w})\varkappa^{i}(\cdot)-\pi_{i}(\cdot)\varkappa^{i}(\tilde{w})=0.$
This proves the first part of the theorem.
To prove the second part, assume that $i_{u}d\alpha=0$, which implies
$\pi_{i}(u)\varkappa^{i}(\cdot)-\pi_{i}(\cdot)\varkappa^{i}(u)=0.$
Note that the 1-forms $\pi_{i},\varkappa^{i}$ for $i=1,\ldots,m$ are linearly independent; therefore,
$\varkappa^{i}(u)=0\quad\text{and}\quad\pi_{i}(u)=0.$
Setting $u=(u_{t},u_{x},u_{p})$, we obtain the relations
$u_{p_{i}}+\frac{\partial H}{\partial x^{i}}u_{t}=0,\quad u_{x^{i}}-\frac{\partial H}{\partial p_{i}}u_{t}=0.$
Thus, $u=u_{t}\tilde{w}$, which completes the proof.
Theorem 11.
The form $\alpha$ is a relative integral invariant of the system (19):
$L_{\tilde{w}}\alpha=d\mathscr{F},\quad\text{where}\quad\mathscr{F}=p_{i}\frac{\partial H}{\partial p_{i}}-H.$
Proof
This can be verified by direct calculation using formula (4). However, it is more convenient to use the Cartan formula and Theorem 10:
$L_{\tilde{w}}\alpha=d(i_{\tilde{w}}\alpha)+i_{\tilde{w}}d\alpha,\quad\text{where}\quad d(i_{\tilde{w}}\alpha)=d\mathscr{F}.$
The theorem is proved.
Note that if the Hamiltonian $H$ admits the Legendre transform with respect to the momenta $p_{i}$, then $\mathscr{F}$ coincides with the Lagrangian function $L(t,x,\dot{x})$ evaluated along the trajectories of the system.
Theorem 12.
Let $\gamma\subset\tilde{N}$ be a closed curve. Then the integral
$\int_{g^{\tau}(\gamma)}\alpha$
is independent of $\tau$.
This fact is a direct consequence of Theorem 3.
Remark 2 implies that the form $d\alpha$ is an integral invariant: $L_{\tilde{w}}d\alpha=0.$
Theorem 13.
Let $\Sigma\subset\tilde{N}$ be an admissible two-dimensional surface. Then the integral
$\int_{g^{\tau}(\Sigma)}d\alpha$
is independent of $\tau$.
This fact is a direct consequence from Theorem 2.
Theorem 14.
Let $\Sigma\subset N$ be a two-dimensional surface. Then the integral
$\int_{G^{t}_{t_{0}}(\Sigma)}\beta$
is independent of $t$ and $t_{0}$.
Proof
This fact is a direct consequence of Theorem 6. Indeed, since $d_{z}\beta=0$ and by virtue of equation (18), we obtain:
$\frac{\partial\beta}{\partial t}+L_{w}\beta=d_{z}i_{w}\beta+i_{w}d_{z}\beta=0.$
The theorem is proved.
Corollary 2.
From theorem 14 it follows that $G^{t}_{t_{0}}$ preserves the form $\beta:$
$(G^{t}_{t_{0}})_{*}\beta=\beta.$ (20)
In other words, the shift along the trajectories of a Hamiltonian system is a symplectic map.
Moreover, equation (20) implies that this shift preserves the volume in the phase space:
$(G^{t}_{t_{0}})_{*}\varpi=\varpi,\quad\varpi=\underbrace{\beta\wedge\ldots\wedge\beta}_{m\text{ times}}.$
Theorem 15.
If $\gamma\subset N$ is a closed curve, then the integral
$\int_{G^{t}_{t_{0}}(\gamma)}\zeta,\quad\text{where}\quad\zeta=p_{i}dx^{i},$
is independent of $t$ and $t_{0}$.
This fact is a direct consequence of Theorem 6. Indeed, using the Cartan formula, we have:
$\frac{\partial\zeta}{\partial t}+L_{w}\zeta=d_{z}i_{w}\zeta+i_{w}d_{z}\zeta=d_{z}\left(p_{i}\frac{\partial H}{\partial p_{i}}-H\right).$
Problem 3.
Consider a Hamiltonian system with a Hamiltonian function $H(t,x,p)$, where $x=(x^{1},\ldots,x^{m})$, $p=(p_{1},\ldots,p_{m})\in\mathbb{R}^{m}$, and $t\in\mathbb{R}$. Assume this Hamiltonian to be positively homogeneous of degree one in the momenta:
$H(t,x,\lambda p)=\lambda H(t,x,p)\quad\forall\lambda>0.$
Recall that such functions satisfy the Euler homogeneous function theorem:
$H=p_{i}\frac{\partial H}{\partial p_{i}}.$
Let $\gamma\subset N$ be an admissible one-dimensional manifold (not necessarily closed). Show that the integral
$\int_{G_{t_{0}}^{t}(\gamma)}p_{i}dx^{i}$
is independent of $t$ and $t_{0}$.
Problem 4.
Let the Hamiltonian $H$ be as defined in Problem 3. Fix the initial position $x(0)=\hat{x}$ and treat the initial momentum $p(0)=\hat{p}$ as a parameter.
[…]
$\frac{\partial x^{i}(t,\hat{p})}{\partial\hat{p}_{s}}p_{i}(t,\hat{p})=0,\quad s=1,\ldots,m.$ (21)
Hint: Verify the equality for $t=0$ and, by direct calculation, show that
$\frac{d}{dt}\left(\frac{\partial x^{i}(t,\hat{p})}{\partial\hat{p}_{s}}p_{i}(t,\hat{p})\right)=0.$
7. Reduction of the Order of a Hamiltonian System by Use of the Energy Integral
Let us point out a consequence of Theorem 10.
Assume that the Hamiltonian $H$ does not depend on $t$. Then $H$ is a first integral. Assume that the energy level
$E_{h}=\{H(z)=h\}\subset\tilde{N}$
is non-degenerate: $dH|_{E_{h}}\neq 0$.
Without loss of generality, we can assume that in any sufficiently small domain of the manifold $E_{h}$, the following inequality holds:
$\frac{\partial H}{\partial p_{1}}\neq 0.$
We can always achieve this by canonical permutations. [^1] Changes of variables $(x,p)\mapsto(X,P)$ of the type $X^{i}=-p_{i},\quad P_{i}=x^{i},\quad P_{s}=p_{s},\quad X^{s}=x^{s},\quad s\neq i$ or $X^{i}=x^{j},\quad X^{j}=x^{i},\quad P_{i}=p_{j},\quad P_{j}=p_{i}$ and compositions of such changes are called canonical permutations. From the results of Section 10, it follows that canonical permutations preserve the Hamiltonian form of the equations.
Then the Implicit Function Theorem implies that the surface $E_{h}$ is described locally by the graph
$p_{1}=g(x^{1},\dots,x^{m},p_{2},\dots,p_{m},h),$ (22)
where $t,x^{1},\dots,x^{m},p_{2},\dots,p_{m}$ are local coordinates on $E_{h}$.
Therefore, we obtain:
$\alpha|_{E_{h}}=\sum_{k=2}^{m}p_{k}dx^{k}-\big(-g(x^{1},\dots,x^{m},p_{2},\dots,p_{m},h)\big)dx^{1}-hdt,\quad d(hdt)=0.$
Thus, the form $d(\alpha|_{E_{h}})$ is annihilated by the vector field associated with the following system:
$\frac{dt}{dT}$ $=b(t,x^{1},\dots,x^{m},p_{1},\dots,p_{m}),$
$\frac{dx^{1}}{dT}$ $=1,$
$\frac{dp_{i}}{dT}$ $=\frac{\partial g}{\partial x^{i}},\quad\frac{dx^{i}}{dT}=-\frac{\partial g}{\partial p_{i}},\quad i=2,\dots,m$ (23)
for an arbitrary function $b$.
Equations (23) have Hamiltonian form. Their trajectories, together with the relation $x^{1}=T+T_{0}$ and equation (22), are the projections of the trajectories of system (19) from $E_{h}$ onto the phase space NN.
8. Characteristic Property of the Hamilton-Jacobi Equation
Theorem 16.
  1. (1)
    Assume that a function $S=S(t,x)$ satisfies the Hamilton-Jacobi equation:
    $H\left(t,x,\frac{\partial S}{\partial x}\right)+\frac{\partial S}{\partial t}=0.$
    Then the graph
    $\Gamma=\left\{p_{i}=\frac{\partial S}{\partial x^{i}}(t,x),\quad i=1,\dots,m\right\}\subset\tilde{N}$
    is an $(m+1)$-dimensional invariant manifold of system (19).
  2. (2)
    Conversely, let a graph
    $\Gamma=\left\{p=\frac{\partial S}{\partial x}\right\}\subset\tilde{N},\quad S=S(t,x)$
    be an invariant surface of system (19). Then there exists a function $\psi=\psi(t)$ such that $S$ satisfies the equation
    $H\left(t,x,\frac{\partial S}{\partial x}\right)+\frac{\partial S}{\partial t}=\psi(t).$
    (Consequently, the function $\tilde{S}=S-\int\psi(t)\,dt$ satisfies the Hamilton-Jacobi equation.)
  3. (3)
    If $S$ is a solution to the Hamilton-Jacobi equation, then $\alpha|_{\Gamma}=dS$.
Proof
The last assertion is trivial. Let us prove item 1).
Observe that $(t,x)$ are local coordinates on the manifold $\Gamma$. Let $x(t)$ be a solution to the system
$\dot{x}^{i}=\frac{\partial H}{\partial p_{i}}\left(t,x,\frac{\partial S}{\partial x}(t,x)\right).$
We shall show that
$x(t),\quad p_{i}(t)=\frac{\partial S}{\partial x^{i}}(t,x(t))$
is a solution to the Hamilton equations.
Indeed, let us introduce the function
$F(t,x)=H\left(t,x,\frac{\partial S}{\partial x}(t,x)\right).$
Then we have
$\frac{\partial F}{\partial x^{i}}=\frac{\partial H}{\partial x^{i}}+\frac{\partial H}{\partial p_{s}}\frac{\partial^{2}S}{\partial x^{i}\partial x^{s}}.$ (24)
Now we calculate $\dot{p}_{k}$:
$\dot{p}_{k}$ $=\frac{d}{dt}\left(\frac{\partial S}{\partial x^{k}}(t,x(t))\right)=\frac{\partial^{2}S}{\partial t\partial x^{k}}+\frac{\partial^{2}S}{\partial x^{k}\partial x^{s}}\dot{x}^{s}$
$=\frac{\partial^{2}S}{\partial t\partial x^{k}}+\frac{\partial^{2}S}{\partial x^{k}\partial x^{s}}\frac{\partial H}{\partial p_{s}}=\frac{\partial^{2}S}{\partial t\partial x^{k}}+\left(\frac{\partial F}{\partial x^{k}}-\frac{\partial H}{\partial x^{k}}\right)$
$=\frac{\partial}{\partial x^{k}}\left(\frac{\partial S}{\partial t}+F\right)-\frac{\partial H}{\partial x^{k}}=-\frac{\partial H}{\partial x^{k}}.$
To prove item 2), take a solution $(x(t),p(t))$ to the Hamilton equations and differentiate the equality
$p_{i}(t)=\frac{\partial S}{\partial x^{i}}(t,x(t))$
with respect to $t$:
$\frac{\partial H}{\partial x^{i}}=\frac{\partial^{2}S}{\partial x^{i}\partial x^{r}}\frac{\partial H}{\partial p_{r}}+\frac{\partial^{2}S}{\partial x^{i}\partial t}.$
Using formula (24), we obtain
$\frac{\partial}{\partial x^{i}}\left(\frac{\partial S}{\partial t}+F\right)=0.$
The theorem is proved.
Item 3) implies two consequences. Firstly, we have $d\alpha|_{\Gamma}=0$. Secondly, if $(\tau,x(\tau),p(\tau))$ is a trajectory of (19) that belongs to the manifold $\Gamma$, then the following equality holds:
$S(t,x(t))-S(t_{0},x(t_{0}))$ $=\int_{t_{0}}^{t}\left(p_{i}(\tau)\dot{x}^{i}(\tau)-H(\tau,x(\tau),p(\tau))\right)d\tau$
$=\int_{t_{0}}^{t}L(\tau,x(\tau),\dot{x}(\tau))\,d\tau,$ (25)
where $L$ is the Lagrangian.
The second equality holds only if the Legendre transform of $H$ is well-defined. Indeed,
$p_{i}(\tau)\dot{x}^{i}(\tau)-H(\tau,x(\tau),p(\tau))=p_{i}(\tau)\frac{\partial H}{\partial p_{i}}-H(\tau,x(\tau),p(\tau)).$
For systems in classical mechanics, there are no issues with the Legendre transform.
Note also that if the Hamiltonian $H$ is a homogeneous function in $p$ (as in Problem 3), then the first line of equality (25) takes the form:
$S(t,x(t))=S(t_{0},x(t_{0})).$
9. Eikonal Equation and Gauss’s Lemma
9.1. Eikonal Equation
Let $M$ be a Riemannian manifold with local coordinates $x=(x^{1},\dots,x^{m})$, and let $g_{ij}(x)$ denote the components of the metric tensor.
Let $x(t)$ be a solution to the Euler-Lagrange equations with the Lagrangian
$L=\frac{1}{2}g_{ij}(x)\dot{x}^{i}\dot{x}^{j},\quad H=\frac{1}{2}g^{ij}(x)p_{i}p_{j},$ (26)
where the initial conditions and conjugate momenta are given by
$x(0)=\hat{x},\quad p_{i}=\frac{\partial L}{\partial\dot{x}^{i}}=g_{ij}(x)\dot{x}^{j}.$
Assume that the solution $x(t)$ has energy $1/2$:
$|\dot{x}(t)|^{2}=g_{ij}(x(t))\dot{x}^{i}(t)\dot{x}^{j}(t)=1.$ (27)
Under this assumption, the curve $x=x(t)$ defines a geodesic, and $t$ is the arc-length parameter.
Theorem 17.
Suppose a function $f\colon M\to\mathbb{R}$ satisfies the eikonal equation
$|\nabla f|^{2}=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial f}{\partial x^{j}}=1,$ (28)
where the gradient is given by $\nabla f=g^{ij}\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}$. Let $x(t)$ be a geodesic issued from a level surface
$\Psi_{\hat{x}}=\{x\in M\mid f(x)=f(\hat{x})\},\quad\hat{x}\in\Psi_{\hat{x}}$,
such that it is orthogonal to this surface at the initial point:
$\frac{\partial f}{\partial x^{i}}(\hat{x})=g_{ij}(\hat{x})\dot{x}^{j}(0).$ (29)
Then the following properties hold:
  1. (1)
    The geodesic remains orthogonal to each level surface $\Psi_{x(t)}$ it intersects:
    $\frac{\partial f}{\partial x^{i}}(x(t))=g_{ij}(x(t))\dot{x}^{j}(t);$
  2. (2)
    The value of the function along the geodesic satisfies:
    $f(x(t))-f(\hat{x})=t.$
Proof of Theorem 17
The function S(t,x)=f(x)t/2S(t,x)=f(x)-t/2 satisfies the Hamilton–Jacobi equation with the Hamiltonian given in (26). Moreover, due to condition (29) and Theorem 16, the geodesic $x(t)$ belongs to the Lagrangian manifold $\Gamma$ defined by:
$p_{i}(t)=g_{ij}(x(t))\dot{x}^{j}(t)=\frac{\partial S}{\partial x^{i}}(t,x(t))=\frac{\partial f}{\partial x^{i}}(x(t)).$
This implies that the velocity vector $\dot{x}(t)$ is orthogonal to the level surface $\Psi_{x(t)}$.
By virtue of formula (27), we have
$\int_{0}^{t}L\,d\tau=\frac{t}{2}.$
Consequently, from the relation for the action (or formula (25)), we find
$S(t,x(t))-S(0,\hat{x})=\frac{t}{2}.$
Substituting S(t,x)=f(x)t/2S(t,x)=f(x)-t/2 into this equality, we obtain
$\left(f(x(t))-\frac{t}{2}\right)-f(\hat{x})=\frac{t}{2},$
which directly implies item 2) of the theorem.
Theorem 17 is proved.
Problem 5.
Suppose a function $f\colon M\to\mathbb{R}$ satisfies the eikonal equation (28). Show that all solutions to the autonomous system of ordinary differential equations
$\dot{x}=\nabla f(x)$ (30)
are geodesics.
Solution to Problem 5
The proof follows the previous argument almost literally.
[…]
$p_{i}(t)=g_{ij}(x(t))\dot{x}^{j}(t)=\frac{\partial S}{\partial x^{i}}(t,x(t))=\frac{\partial f}{\partial x^{i}}(x(t)).$
Consequently, $x(t)$ is a solution to (30).
There is, however, a more concise way to solve Problem 5. Any solution to (30) minimizes the following action functional:
$I[x]=\int_{t_{1}}^{t_{2}}\frac{1}{2}|\dot{x}-\nabla f|^{2}\,dt.$
Note that the integrand can be expanded as
$\frac{1}{2}|\dot{x}-\nabla f|^{2}=\frac{1}{2}|\dot{x}|^{2}+\frac{1}{2}|\nabla f|^{2}-\langle\dot{x},\nabla f\rangle=\frac{1}{2}|\dot{x}|^{2}+\frac{1}{2}-\frac{df}{dt}.$
Since $\frac{1}{2}$ and the total derivative $df/dt$ do not affect the Euler–Lagrange equations, any solution to (30) must also be a solution to the equations of motion for the Lagrangian $L=\frac{1}{2}|\dot{x}|^{2}$.
Problem 6.
Let $f(x,a)$ be an $\ell$-parametric family of solutions to the eikonal equation (28), where $a=(a^{1},\dots,a^{\ell})\in\mathbb{R}^{\ell}$ denotes the parameters.
Assume that the system of equations
$\frac{\partial f}{\partial a}(x,a)=0$ (31)
possesses a smooth solution $a=\theta(x)$.
Show that the envelope defined by
$\tilde{f}(x):=f(x,\theta(x))$
is also a solution to equation (28) (see [5]).
Furthermore, show that for any given value of $a$, the level surface $\{\tilde{f}(x)=c\}$ is tangent to the corresponding surface $\{f(x,a)=c\}$ of the family at their points of intersection.
9.2. Gauss’s Lemma
Consider the result of Problem 4 from a Riemannian geometry perspective.
The Hamiltonian
$\mathcal{H}=\sqrt{g^{ij}(x)p_{i}p_{j}}$
is a homogeneous function of degree one in the momenta. The function $\mathcal{H}$ is a first integral of the system. Let $(x(t),p(t))$ be a solution with a fixed initial position $x(0)=\hat{x}$ and satisfying the condition
$\mathcal{H}=1.$ (32)
We regard the initial momentum $p(0)=\hat{p}$ as a parameter.
It is straightforward to see that such solutions also satisfy the system with the Hamiltonian given in (26). In this case, condition (32) is equivalent to (27). In particular, $t$ is the arc-length parameter, and solutions $x(t)$ of the type described above are geodesics on $M$ emanating from the point $\hat{x}$.
Consider a sphere of radius $r>0$ in the tangent space $T_{\hat{x}}M$:
$C_{r}=\{v=(v^{1},\dots,v^{m})\in T_{\hat{x}}M\mid|v|^{2}=g_{ij}(\hat{x})v^{i}v^{j}=r^{2}\}.$
Each vector $v\in C_{r}$ defines a unique geodesic $x_{v}(t)$ satisfying the initial conditions $x_{v}(0)=\hat{x}$ and $\dot{x}_{v}(0)=v$.
Let $r$ be chosen sufficiently small so that any two distinct geodesics from the family $\{x_{v}\}_{v\in C_{r}}$ do not intersect for $t\in(0,1]$. Under this condition, the set
$\Sigma=\{x_{v}(1)\in M\mid v\in C_{r}\}$
is a smooth submanifold of $M$ of dimension m1m-1.
Theorem 18.
The velocity vector $\dot{x}_{v}(1)\in T_{x_{v}(1)}M$, where $|v|=r$, is orthogonal to the tangent space of the submanifold $\Sigma$ at the point $x_{v}(1)$.
Indeed, consider an arbitrary smooth curve $x(\xi)=x_{v(\xi)}(1)$ on $\Sigma$, where $|v(\xi)|=r$ and $\xi\in\mathbb{R}$ is a parameter along the curve.
A tangent vector to $\Sigma$ at the point $x_{v}(1)$ is given by:
$\frac{dx^{i}_{v(\xi)}(1)}{d\xi}=\frac{\partial x^{i}_{v}}{\partial v^{r}}\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}\frac{\partial\hat{p}_{j}}{\partial v^{r}}\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}g_{jr}(\hat{x})\frac{dv^{r}}{d\xi}=\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}\frac{d\hat{p}_{j}}{d\xi}.$
By virtue of (21), the following orthogonality condition holds:
$g_{il}(x_{v}(1))\dot{x}^{l}_{v}(1)\frac{\partial x^{i}_{v}}{\partial\hat{p}_{j}}=0.$
Consequently, it follows that
$g_{il}(x_{v}(1))\dot{x}^{l}_{v}(1)\frac{d}{d\xi}x^{i}_{v}(1)=0,$
which means the velocity vector $\dot{x}_{v}(1)$ is orthogonal to every tangent vector of $\Sigma$. This completes the proof of the theorem.
10. Canonical Transformations. Generating Functions
10.1. Canonical Transformations
Let $f=f(t,x,p)$ be a function defined on the extended phase space $\tilde{N}$ of the system (16). Recall the notation:
$d_{z}f=\frac{\partial f}{\partial x^{i}}dx^{i}+\frac{\partial f}{\partial p_{i}}dp_{i},\quad df=d_{z}f+\frac{\partial f}{\partial t}dt.$
Definition 3 ([1]).
A transformation [^2] We use the word “transformation” as a synonym for the phrase “change of variables.” of the manifold $\tilde{N}$
$(t,x,p)\mapsto(t,X,P),\quad X=X(t,x,p),\quad P=P(t,x,p)$
is said to be canonical if
$d_{z}P_{i}\wedge d_{z}X^{i}=dp_{i}\wedge dx^{i}.$ (33)
In the following, we treat $t,x,p$ as independent variables, such that
$d_{z}p_{i}=dp_{i},\quad d_{z}x^{i}=dx^{i},\quad d_{z}t=0.$
Loosely speaking, a canonical transformation is a family of symplectic maps parameterized by $t$. For instance, according to Corollary 2, the mapping $(t,x,p)\mapsto(t,G^{t}_{t_{0}}(x,p))$ is a canonical transformation.
Equation (33) is locally equivalent to the existence of a function $S(t,x,p)$ such that
$p_{i}dx^{i}-P_{i}d_{z}X^{i}=d_{z}S(t,x,p).$ (34)
Indeed, this follows from the fact that the exterior derivative vanishes: $d_{z}(p_{i}dx^{i}-P_{i}d_{z}X^{i})=0$.
Equality (34) can be rewritten in the form
$P_{i}dX^{i}+P_{i}\frac{\partial X^{i}}{\partial t}dt+p_{i}dx^{i}=dS-\frac{\partial S}{\partial t}dt.$ (35)
Theorem 19.
In the new coordinates $(X,P)$, the Hamiltonian system (16) preserves its canonical form:
$\dot{P}_{i}=-\frac{\partial K}{\partial X^{i}},\quad\dot{X}^{i}=\frac{\partial K}{\partial P_{i}},$
where the new Hamiltonian $K(t,X,P)$ is given by
$K(t,X,P)=\left(P_{i}\frac{\partial X^{i}}{\partial t}+\frac{\partial S}{\partial t}+H\right)\bigg|_{(x,p)\to(X,P)}.$
Indeed, formula (35) implies that
$\alpha=p_{i}dx^{i}-Hdt=P_{i}dX^{i}-Kdt+dS,$ (36)
and the assertion of Theorem 19 follows directly from Theorem 10.
Corollary 3.
If a canonical transformation does not depend on time,
$P=P(x,p),\quad X=X(x,p),$
then the Hamiltonian $H$ transforms as a scalar function on the manifold $N$:
$K(t,X,P)=H\big(t,x(X,P),p(X,P)\big).$ (37)
10.2. Generating Functions
Definition 4.
A canonical transformation $(t,x,p)\mapsto(t,X,P)$ is said to be free if
$\det\left(\frac{\partial X^{i}}{\partial p_{j}}\right)\neq 0.$ (38)
In this case, by the Implicit Function Theorem, the variables $(t,x,X)$ serve as local coordinates on $\tilde{N}$. In particular, we can express the generating function as $S=S_{1}(t,x,X)$.
Formula (36) then takes the form
$p_{i}dx^{i}-Hdt=P_{i}dX^{i}-Kdt+\frac{\partial S_{1}}{\partial t}dt+\frac{\partial S_{1}}{\partial x^{i}}dx^{i}+\frac{\partial S_{1}}{\partial X^{i}}dX^{i},$ (39)
which implies the following relations:
$p_{i}$ $=\frac{\partial S_{1}}{\partial x^{i}},\quad P_{i}=-\frac{\partial S_{1}}{\partial X^{i}};$ (40)
$K$ $=H+\frac{\partial S_{1}}{\partial t}.$ (41)
The condition (38) can be written as
$\det\left(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\right)\neq 0.$ (42)
Conversely, if the inequality (42) holds, then the relations (40) define a canonical transformation.
Indeed, let us verify the non-degeneracy of the mapping $(x,p)\mapsto(X,P)$. We can represent this transformation as a composition of two steps:
$(x,p)\mapsto(x^{\prime},X^{\prime})\mapsto(X,P).$
The first step is the inverse of the transformation defined by:
$x=x^{\prime},\quad p=\frac{\partial S_{1}}{\partial x}(t,x^{\prime},X^{\prime}).$
The Jacobian of this change of variables is
$\det\left(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\right)^{-1}.$
The second step is given by the formulas:
$X=X^{\prime},\quad P=-\frac{\partial S_{1}}{\partial X}(t,x^{\prime},X^{\prime}).$
Its Jacobian is
$\det\left(\frac{\partial^{2}S_{1}}{\partial x^{i}\partial X^{j}}\right).$
Since the total Jacobian (the product of these two) is non-zero, the transformation is well-defined. Therefore, if the function $S_{1}$ satisfies (42), it defines a free transformation which is uniquely determined by the relations (40).
The function $S_{1}$ is called a generating function.
In the new variables $(X,P)$, the Hamiltonian $K$ is given by (41).
Remark 4.
Assume that we have a generating function $S_{1}(t,x,X)$ that satisfies the Hamilton–Jacobi equation:
$H\left(t,x,\frac{\partial S_{1}}{\partial x}\right)+\frac{\partial S_{1}}{\partial t}=0.$
Then, the transformation $(t,x,p)\mapsto(t,X,P)$ is determined, and in the new variables $(X,P)$, the new Hamiltonian $K$ vanishes identically ($K\equiv 0$).
Consequently, the equations of motion in the new coordinates take the form $\dot{X}=0$ and $\dot{P}=0$, meaning the system is integrated immediately.
Consider a composition of canonical transformations:
$(x,p)\mapsto(\tilde{X},\tilde{P})\mapsto(X,P).$
[…]
$p_{i}=\frac{\partial S_{1}}{\partial x^{i}},\quad\tilde{P}_{i}=-\frac{\partial S_{1}}{\partial\tilde{X}^{i}}.$ (43)
The second transformation is a canonical permutation:
$\tilde{X}^{i}=P_{i},\quad\tilde{P}_{i}=-X^{i}.$
Thus, the change $(x,p)\mapsto(X,P)$ is expressed by means of a generating function $S_{2}(t,x,P):=S_{1}(t,x,P)$. Formulas (43) then imply:
$p_{i}=\frac{\partial S_{2}}{\partial x^{i}},\quad X^{i}=\frac{\partial S_{2}}{\partial P_{i}}.$
In this case, the non-degeneracy condition (42) takes the form
$\det\left(\frac{\partial^{2}S_{2}}{\partial x^{i}\partial P_{j}}\right)\neq 0,$
and the condition (38) for a free transformation is written as
$\det\left(\frac{\partial P_{i}}{\partial p_{j}}\right)\neq 0.$
By applying canonical permutations to various subsets of the conjugate pairs $(\tilde{X}^{i},\tilde{P}_{i})$, we can obtain $2^{m}$ different types of generating functions.
Remark 5.
The identity mapping
$P=p,\quad X=x$
is obtained via the $S_{2}$-type generating function $S_{2}=P_{i}x^{i}$.
Definition 5.
A function $S=S(t,x,b)$ depending on $m$ parameters $b=(b_{1},\dots,b_{m})$ is called a complete integral of the Hamilton–Jacobi equation
$H\left(t,x,\frac{\partial S}{\partial x}\right)+\frac{\partial S}{\partial t}=0,$
provided that it satisfies the equation for all admissible values of $b$ and the non-degeneracy condition
$\det\left(\frac{\partial^{2}S}{\partial x^{i}\partial b_{j}}\right)\neq 0$
holds.
Therefore, if a complete integral is provided, the system of Hamilton’s equations can be integrated in closed form. See Remark 4.
The vector $b$ can be interpreted as a set of new coordinates, in which case we obtain the generating function $S_{1}$. Alternatively, if we treat $b$ as a set of new momenta, we obtain $S_{2}$. More generally, the vector $b$ can consist of a mixture of both new coordinates and new momenta.
Currently, most known integrable Hamiltonian systems are integrated using the separation of variables procedure. For the autonomous case ($H=H(z)$), this means that in certain suitable coordinates, the Hamilton–Jacobi equation
$H\left(x,\frac{\partial S}{\partial x}\right)=K(b)$
admits a complete integral of the form
$S(x,b)=\sum_{k=1}^{m}S_{k}(x^{k},b).$
11. Hamiltonian Vector Field Straightening Theorem
Assume that the system (16) is autonomous, i.e., $H=H(z)$.
Theorem 20.
Assume that the Hamiltonian $H$ is non-degenerate at a point $\tilde{z}\in N$:
$dH(\tilde{z})\neq 0.$
Then, in some neighborhood $U$ of the point $\tilde{z}$, there exist canonical coordinates
$Z=(X,P)=(X^{1},\dots,X^{m},P_{1},\dots,P_{m}),\quad Z=Z(x,p)$
such that:
  1. (1)
    $dp_{i}\wedge dx^{i}=dP_{i}\wedge dX^{i}$ (the transformation is canonical);
  2. (2)
    In the coordinates $Z$, the Hamiltonian takes the form $H=X^{1}$.
Proof
Without loss of generality, let $\tilde{z}=0$ and
$H(0)=0,\quad\frac{\partial H}{\partial p_{1}}(0)\neq 0.$ (44)
Then, by the Implicit Function Theorem, the equation
$H(x,p)=X^{1}$ (45)
has a solution
$p_{1}=\phi(x,p_{2},\dots,p_{m},X^{1}),\quad H(x,\phi,p_{2},\dots,p_{m})=X^{1}$
for sufficiently small $|X^{1}|$, $|x|$, and $|p|$ (cf. (22)). From (44), 45 we obtain
$\phi(0)=0,\quad\frac{\partial H}{\partial p_{1}}(0)\cdot\frac{\partial\phi}{\partial X^{1}}(0)=1.$ (46)
Consider the following Cauchy problem:
$\frac{\partial S}{\partial x^{1}}=\phi\left(x,\frac{\partial S}{\partial x^{2}},\dots,\frac{\partial S}{\partial x^{m}},X^{1}\right),\quad S\big|_{x^{1}=0}=\sum_{k=2}^{m}x^{k}X^{k},\quad S=S(x,X).$ (47)
As shown in Section 13, this Cauchy problem has a solution $S=S(x,X)$ for sufficiently small $|x|$ and $|X|$.
Thus, equation (47) is equivalent to the equation
$H\left(x,\frac{\partial S}{\partial x}\right)=X^{1}.$
To complete the proof, it remains to verify that $S(x,X)$ is a generating function of a transformation $(x,p)\mapsto(X,P)$:
$p=\frac{\partial S}{\partial x},\quad P=-\frac{\partial S}{\partial X}.$
The values $x=0,p=0$ correspond to $X=0,P=0$.
By virtue of (47), we have
$\frac{\partial^{2}S}{\partial x\partial X}\bigg|_{x=X=0}=\begin{bmatrix}\frac{\partial\phi}{\partial X^{1}}&\frac{\partial^{2}S}{\partial x^{1}\partial X^{2}}&\dots&\frac{\partial^{2}S}{\partial x^{1}\partial X^{m}}\\ 0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&1\end{bmatrix},\quad\det\frac{\partial^{2}S}{\partial x\partial X}\bigg|_{x=X=0}=\frac{\partial\phi}{\partial X^{1}}.$
From (46), it follows that
$\frac{\partial\phi}{\partial X^{1}}(0)\neq 0.$
The theorem is proved.
12. The Poincare Section on the Energy Level, [10]
Recall that the Hamiltonian vector field $w$ is defined by formula (17). We assume that $H$ does not depend on $t$.
Let $E_{h}=\{z\in N\mid H(z)=h\}$ denote an energy level, assuming $dH|_{E_{h}}\neq 0$.
Let $Y\subset E_{h}$ be a hypersurface of dimension $\dim Y=2m-2$ such that $w(z)\notin T_{z}Y$ for all $z\in Y$. Consider the restriction $\beta^{\prime}=\beta|_{Y}$.
Theorem 21.
The form $\beta^{\prime}$ is nondegenerate; thus, $(Y,\beta^{\prime})$ is a symplectic manifold.
Proof
Take a point $z\in Y$. In a neighborhood of this point, introduce canonical coordinates in accordance with Theorem 20. In these coordinates, the integral curves are given by:
$P_{1}(t)=-t+P_{1}(0),\quad P_{j}(t)=P_{j}(0),\quad X^{i}(t)=X^{i}(0),\quad j=2,\dots,m,$
and the energy level is $E_{h}=\{X^{1}=h\}$.
Thus, $(X^{2},\dots,X^{m},P_{1},\dots,P_{m})$ serve as local coordinates on $E_{h}$, and the manifold $Y$ can be represented as a graph $P_{1}=y(X^{2},\dots,X^{m},P_{2},\dots,P_{m})$. This implies that $(X^{2},\dots,X^{m},P_{2},\dots,P_{m})$ are local coordinates on $Y$.
The proof is completed by the following computation:
$\beta|_{Y}=\left.\sum_{i=1}^{m}dP_{i}\wedge dX^{i}\right|_{Y}=\sum_{k=2}^{m}dP_{k}\wedge dX^{k}.$
The theorem is proved.
Let $G^{t}:N\to N$ denote the phase flow of (16). Recall that $G^{t}_{*}\beta=\beta$. Assume that the manifolds $Y_{1},Y_{2}\subset E_{h}$ satisfy the same properties as $Y$. As proved above, the pairs $(Y_{i},\beta_{i})$, where $\beta_{i}=\beta|_{Y_{i}}$, are symplectic manifolds.
Suppose that a trajectory starting at $z\in Y_{1}$ reaches the surface $Y_{2}$ at time $t=\tau(z)>0$, such that:
$Q(z):=G^{\tau(z)}(z)\in Y_{2}.$
Theorem 22.
The mapping $Q\colon Y_{1}\to Y_{2}$ is symplectic; that is, $Q_{*}\beta_{2}=\beta_{1}$.
Proof
Take a two-dimensional compact manifold $\Sigma\subset Y_{1}$ with a smooth boundary curve $\partial\Sigma$ without self-intersections and consider a cylinder
$C=\bigcup_{z\in\Sigma}\{G^{t}(z)\mid t\in[0,\tau(z)]\}.$
Since the form $\beta$ is closed, we have
$\int_{\partial C}\beta=0.$
On the other hand, it follows that
$\partial C=\Sigma\cup Q(\Sigma)\cup\sigma,\quad\sigma=\bigcup_{z\in\partial\Sigma}\{G^{t}(z)\mid t\in(0,\tau(z))\}.$
By a suitable choice of orientations, we obtain
$\int_{\Sigma}\beta=\int_{Q(\Sigma)}\beta+\int_{\sigma}\beta.$
Since $\Sigma$ is an arbitrary compact manifold, to complete the proof, it is enough to show that
$\beta|_{\sigma}=0.$
A basis in $T_{z}\sigma$ can be chosen as follows: $w,v$, where $w$ is the Hamiltonian vector field and $v$ is some other vector.
Let us calculate $\beta$ on the basis vectors by using formula (18):
$\beta(w,v)=-dH(v)=0.$
The last equality holds because $v\in T_{z}\sigma\subset T_{z}E_{h}$.
The theorem is proved.
The result of Theorem 22 remains valid if $Y_{1}=Y_{2}$.
13. The Hamilton–Jacobi Equation in the General Case: The Method of Characteristics
In the theory of PDEs, the equation
$u_{t}+f\left(t,x,u,\frac{\partial u}{\partial x}\right)=0$ (48)
is also called the Hamilton–Jacobi equation.
Here, $f=f(t,x,\xi,p)$ is a scalar function of the variables
$t,\xi\in\mathbb{R},\quad p=(p_{1},\dots,p_{m}),\quad x=(x^{1},\dots,x^{m})\in\mathbb{R}^{m}.$
Consider the following system of ODEs:
$\dot{\xi}$ $=p_{i}\frac{\partial f}{\partial p_{i}}-f,$ (49)
$\dot{p}_{i}$ $=-\frac{\partial f}{\partial x^{i}}-\frac{\partial f}{\partial\xi}p_{i},$
$\dot{x}^{i}$ $=\frac{\partial f}{\partial p_{i}}.$
The extended phase space $V$ of this system is the space of variables
$V=\{(t,x,\xi,p)\in\mathbb{R}^{2m+2}\}.$
If the function $f$ does not depend on $\xi$, then the equations for $x$ and $p$ in (49) decouple and form a system of Hamiltonian equations, while equation (48) becomes a Hamilton–Jacobi equation in the sense discussed above.
System (49) is called the characteristic system, and its solutions are called characteristics.
The proof of the following theorem is analogous to the proof of Theorem 16.
Theorem 23.
Let $u=u(t,x)$ be a solution to (48). Then the $(m+1)$-dimensional manifold (graph)
$G=\left\{\xi=u(t,x),\quad p_{i}=\frac{\partial u}{\partial x^{i}}(t,x),\quad i=1,\dots,m\right\}\subset V$
is an invariant manifold of system (49), and the following formula holds:
$(p_{i}dx^{i}-fdt-d\xi)\big|_{G}=0,$
or equivalently,
$(p_{i}dx^{i}-fdt)\big|_{G}=du.$
This theorem provides a method for solving the Cauchy problem for equation (48). Indeed, consider this equation with the following initial data:
$u\big|_{t=0}=\hat{u}(x).$ (50)
Define the initial conditions for system (49) as follows:
[…]
Then, the solution $u(t,x)$ to the Cauchy problem (48), (50) is determined by the equations
[…]
By the Implicit Function Theorem, the second equation can be solved for $\hat{x}$ provided that $|t|$ is sufficiently small.
This approach to solving the Cauchy problem (48), (50) is known as the method of characteristics.